### v1

# Asymptotic confidence bands for the histogram regression estimator

Abstract
In a multivariate nonparametric regression model with a Hölder continuous regression function and heteroscedastic noise asymptotic uniform confidence bands are constructed based on the histogram estimator. The radius of the confidence bands does not depend on an extreme value distribution, but instead can be numerically calculated for the chosen binning.
Keywords: Multivariate nonparametric regression, histogram, confidence bands, heteroscedastic noise
MSC 2020: 62G08, 62G20, 62H12
1 Introduction
We consider classical histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise, and derive asymptotic confidence bands.
There is a substantial amount of literature on confidence bands based on different nonparametric regression estimators. Johnston, (1982) has constructed confidence bands for the Nadaraya-Watson kernel estimator. Härdle, (1989) has shown asymptotic uniform confidence bands for a wider class of regression estimators, the M-smoothers. Typically, confidence bands rely on an undersmoothing of the estimator to reduce the bias relative to the stochastic error. An alternative is a direct bias correction, which is used, for instance, by Eubank and Speckman, (1993), who have considered a deterministic, uniform design for local constant regression estimation, and by Xia, (1998) considering a random design under dependence and local linear estimation. There are also bootstrap confidence bands for nonparametric regression, see for instance Hall, (1993), Neumann and Polzehl, (1998), Claeskens and Van Keilegom, (2003) or Hall and Horowitz, (2013). Sabbah, (2014) has shown confidence bands for quantile estimators which is an alternative to a regression model, and Birke et al., (2010) have proved confidence bands in an inverse regression model.
The aforementioned articles only considered univariate regression. In the multivariate case Konakov and Piterbarg, (1984) have analyzed the asymptotic distribution of the maximal deviation for the Nadaraya-Watson estimate in a random design setting. Proksch, (2016) has proved confidence bands in a multivariate regression model with fixed, deterministic design for a general class of estimators that includes, for example, local polynomial estimators. Chao et al., (2017) have covered confidence bands for multivariate quantile and expectile regression with mean regression as a special case.
Despite the popularity and widespread application of histogram estimators, the only uniform confidence band result for a histogram estimator that we are aware of is by Smirnov, (1950) and covers univariate density estimation. We want to close this gap by presenting a confidence band result in a multivariate regression model.
In an undersmooting regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$. The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,J}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $J$ is given by the number of cells in the partition of the histogram. In particular, our confidence bands do not rely on an extreme value distribution. We also incorporate the estimation of the unknown variance and a possible unknown distribution of the covariates.
We obtain a confidence band result for the simple case of a Hölder continuous regression function for every covariate dimension, without a continuity assumption on the covariate density. In contrast for smooth nonparametric regression estimators typically it is assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density have to increase, see Chao et al., (2017), for instance. Instead we apply error moment existence, but the assumptions are not restrictive in the sense of exponential moments.
A central part of the proof for the histogram based confidence band is the application of a result by Chernozhukov et al., 2014b regarding the Gaussian approximation for the suprema of empirical processes. Using this direct approximation of the supremum, it is not necessary to approximate a whole empirical process uniformly. This method has also been applied by Patschkowski and Rohde, (2019) to construct adaptive confidence bands for probability densities. It is an important difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
The main result is presented in Section 2. The proof strategy is outlined in Section 3, while detailed proofs are postponed to Section 4.
2 Histogram estimator and confidence bands
On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in[0,1]^{p}$ and $Y\in\mathbb{R}$. The goal is to estimate the regression function $m\colon[0,1]^{p}\to\mathbb{R}$ with
$Y=m(X)+\varepsilon$
for an observation error $\varepsilon$ with $\mathbb{E}[\varepsilon|X]=0$ and variance function $\sigma^{2}(x)=\mathrm{Var}(\varepsilon|X=x)$. We further assume that the density $f_{X}$ of the covariate or feature vector $X$ and the variance function are bounded from below and above, i.e.
$0<c_{X}\leq f_{X}(\cdot)\leq C_{X}$ (1)
$0<c_{\sigma^{2}}\leq\sigma^{2}(\cdot)\leq C_{\sigma^{2}}$ (2)
for some constants $c_{X},C_{X},c_{\sigma^{2}},C_{\sigma^{2}}>0$.
The histogram estimator relies on a local constant approximation of $m$. For a given partition of the feature space $0,1]^{p}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition. We will choose an equidistant grid to construct the partition such that all cells are hypercubes of the same size. For $\delta^{-1}\in\mathbb{N}$ we consider the grid with edge length $\delta$ on $0,1]^{p}$. Here $\delta=\delta_{n}$ is a sequence depending on the sample size with $\delta\to 0$ for $n\to\infty$. For $x\in[0,1]^{p}$ let $A_{\delta}(x)$ be the hypercube with volume $\delta^{p}$ that contains $x$. Note that the equidistant grid is considered for simplicity and all results carry over to more irregular partitions as long as the volume and the diameters of all cells in the same order. The histogram estimator is then defined by
$\hat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
where $\mathbb{I}\{\cdot\}$ denotes the indicator function. If there is no observation in $A_{\delta}(x)$ we define the estimator as zero on this cell.
In the following theorem we present a confidence band result based on the histogram estimator. Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $0,1]^{p}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
Theorem 2.1.
Let $\alpha\in(0,1],C_{H}>0$ and assume that
$n\delta^{2\alpha+p}(\log n)^{2}\to 0.$ (3)
Further let $4\leq\nu\in\mathbb{N}$ be such that $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ and
$\frac{(\log n)^{5}}{\delta^{p}n^{1-2/\nu}}$ $\to 0$ (4)
and assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$. For independent standard normally distributed random variables $(Z_{j})_{j\in\mathbb{N}}$ and some $\beta\in(0,1)$ let $c_{\delta}(\beta)>0$ be given by
$\mathbb{P}\left(\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\leq c_{\delta}(\beta)\right)=1-\beta.$
For
$\tau_{\delta}(x)=\frac{\mathbb{P}(X\in A_{\delta}(x))^{2}}{\mathbb{E}[\sigma^{2}(X)\mathbb{I}_{A_{\delta}(x)}(X)]}$ (5)
the confidence band
$\mathcal{C}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2}\right]$
satisfies
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
The proof strategy is explained in Section 3, and the proofs are given in Section 4. While the theorem assumes the covariate distribution and variance function to be known, we will consider a replacement of $\tau_{\delta}(x)$ by an estimator below. Note that by Gaussian concentration, we have
$\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2p\log\delta^{-1}})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (6)
due to $\log\delta^{-1}\lesssim\log n$ which is implied by assumption (4). Therefore, $c_{\delta}(\beta)\lesssim\sqrt{\log n}$.
Define
$p_{\delta}(x)=\mathbb{P}(X\in A_{\delta}(x))=\int_{A_{\delta}(x)}f_{X}(t)\,dt.$
By (1) and (2) we have
$c_{X}\delta^{p}\leq p_{\delta}(x)\leq C_{X}\delta^{p}$ (7)
and
$\frac{c_{X}}{C_{\sigma^{2}}}\delta^{p}\leq\tau_{\delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\delta^{p}.$ (8)
Therefore, the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)/(n\delta^{p})}$ uniformly in $x\in[0,1]^{p}$. In the homoscedastic case with constant $\sigma^{2}(\cdot)\equiv\sigma^{2}$ we obtain $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$. For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\delta}(x)=\delta^{p}$ and in that case the confidence band simplifies to
$\mathcal{C}_{n}(x)=\Big{[}\hat{m}(x)-c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}},\hat{m}(x)+c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}}\Big{]}.$
The assumptions (3) and (4) in Theorem 2.1 are up to logarithms that
$n\delta^{p+2\alpha}\to 0\text{ and }n\delta^{p}/(n^{2/\nu})\to\infty,$
where $2/\nu$ is arbitrary small if all moments of $\varepsilon$ exist. Since the approximation error of the histogram estimator is of order $\delta^{\alpha}$ while the stochastic error term is of order $(n\delta^{p})^{-1/2}$, the first assumption is an undersmoothing condition. The second condition implies that the average number of observations in all cells goes to infinity such that the estimator is consistent. As $p$ increases, the assumptions become more restrictive which is the common curse of dimensionality problem for nonparametric function estimation. However, the assumptions are not contradictory, and for every dimension $p$ no bias reduction is needed. In contrast, the confidence bands by Chao et al., (2017) require increasing smoothness orders for the regression function $m$ and covariate density $f_{X}$ if the dimension $p$ increases.
We are not aware of any other confidence band result in the literature for the histogram estimator in the multivariate or univariate regression case. However, the structure of the histogram with fixed deterministic cells might imply that the difference in the proof from the univariate to the multivariate case is minor. The first result for the histogram by Smirnov, (1950) studies univariate density estimation. In addition to the different estimation problem, this result differs from our result in another crucial aspect. The asymptotic bands rely on an extreme value distribution that is independent of $\delta$. Instead our result provides bands with explicit dependence on $\delta$. Thus, the knowledge of the limit distribution is not necessary, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution. Applying an approximation by the maximum of Gaussian random variables instead of a limit distribution is also suggested by Kreiß et al., (2025) for confidence bands for the spectral density in time series analysis.
From Theorem 2.1 we obtain a uniform rate of convergence of the histogram estimator.
Corollary 2.2.
Assume $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}/(\delta^{p}n^{1-2/\nu})\to 0$. Then we have
$\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}}\right).$
For $\nu>2+p/\alpha$ and $\delta=(n/\log n)^{-1/(2\alpha+p)}$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
This convergence rate coincides the minimax-optimal rate for the uniform loss, cf. Tsybakov, (2009). The proof is given in Section 4.
We will now consider the case where the covariate distribution and the variance function of the noise are unknown. To this end, let $\hat{\tau}_{\delta}(x)$ be an estimator for $\tau_{\delta}(x)$ from (5) and define the confidence band as
$\hat{\mathcal{C}}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2}\right].$
Corollary 2.3.
Under the assumptions of Theorem 2.1 and
$\sup_{x\in[0,1]^{p}}\left|\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2})$
we have
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\hat{\mathcal{C}}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
The proof is again postponed to Section 4. To construct the estimator $\hat{\tau}_{\delta}(x)$, we first consider the homoscedastic case where $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$ can be estimated by $\hat{\tau}_{\delta}(x)=\hat{p}_{\delta}(x)/\hat{\sigma}^{2}$ with
$\hat{p}_{\delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}.$ (9)
Note that $\hat{p}_{\delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution. This estimator has the desired rate which we show in Lemma 4.2. The constant variance can be estimated using residuals based on the histogram regression estimator, i.e.
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-\hat{m}(X_{i})\big{)}^{2}.$
Other variance estimators under homoscedasticity and variance function estimators under heteroscedasticity are considered by Shen et al., (2020), among others. Nevertheless, one can also apply a histogram estimator for the variance function $\sigma^{2}(x)=\mathbb{E}[Y^{2}|X=x]-m(x)^{2}$ defined by
$\hat{\sigma}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}-\hat{m}(x)^{2}$
$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\hat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(((\log n)/(n\delta^{p}))^{1/2})$ holds as in Corollary 2.2 under the same assumptions for observations $(X,Y^{2})$, in particular moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$. The histogram estimator $\hat{\sigma}^{2}(\cdot)$ is constant on each hypercube $A_{\delta}(x)$ and thus
$s_{\delta}(x)=\mathbb{E}\big{[}\sigma^{2}(X)\mathbb{I}_{A_{\delta}(x)}(X)\big{]}=\int_{A_{\delta}(x)}\sigma^{2}(t)f_{X}(t)\,dt$ (10)
is approximated by $\int_{A_{\delta}(x)}\hat{\sigma}^{2}(x)f_{X}(t)\,dt=\hat{\sigma}^{2}(x)p_{\delta}(x)$. Hence, we can estimate $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$ by
$\hat{\tau}_{\delta}(x)=\frac{\hat{p}_{\delta}(x)}{\hat{\sigma}^{2}(x)}$
and obtain the rate assumed in Corollary 2.3.
The considered classical histogram estimator is a relatively simple estimator, but it has some similarities to random forests, in particular, the piecewise constant approximation. In a current work we apply a similar proof strategy as explained in the next section to derive confidence bands for centered purely random forests. Presumably, this proof strategy can also be used to obtain confidence bands for other nonparametric regression estimators.
3 Proof strategy
In this section we outline the proof. The estimator error can be decomposed as
$\hat{m}(x)-m(x)=\hat{m}^{(m)}(x)-m(x)+\hat{m}^{(\varepsilon)}(x)$ (11)
with approximation error $\hat{m}^{(m)}(x)-m(x)$ and stochastic error term $\hat{m}^{(\varepsilon)}(x)$, where
$\hat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\qquad\text{and}$
$\hat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
First, we discuss a uniform bound on the approximation error. Note that $\hat{m}^{(m)}(x)=\mathbb{E}\left[\hat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise. Therefore, the approximation error is indeed caused by the ability of the piecewise constant estimator to approximate $m$. Moreover, note that the expectation of the approximation error is the bias of the estimator. For any $A\subset\mathbb{R}^{p}$ let us denote its diameter
$\mathfrak{d}(A):=\sup_{x_{1},x_{2}\in A}\|x_{1}-x_{2}\|,$
where $\|\cdot\|$ is the Euclidean norm. Suppose that there exists at least one observation in $A_{\delta}(x)$. Then we apply that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in[0,1]^{p}$ that
$|\hat{m}^{(m)}(x)-m(x)|$ $=\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}$
$\leq C_{H}\sum_{j=1}^{n}\|X_{j}-x\|^{\alpha}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
$\leq C_{H}\mathfrak{d}(A_{\delta}(x))^{\alpha}=C_{H}(\sqrt{p}\delta)^{\alpha},$ (12)
where we used $\mathfrak{d}(A_{\delta}(x))=\delta\sqrt{p}$ by the Euclidean geometry of the cells. In particular, the approximation error profits from a smaller diameter or mesh size $\delta$, respectively, of the grid cells.
To handle the stochastic error, we can replace the denominator in $\hat{m}^{(\varepsilon)}(x)$ by its expectation, such that
$\tilde{m}^{(\varepsilon)}(x):=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}p_{\delta}(x)^{-1}\mathbb{I}\{X_{j}\in A_{\delta}(x)\}$ (13)
becomes the asymptotically leading term of the stochastic error $\hat{m}^{(\varepsilon)}(x)$. Handling the remainder term $\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 4.3.
The idea for the leading term is that $\tilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough. We write $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$, see (5) and (10). Note that $p_{\delta}(x)$, $\tau_{\delta}(x)$ and $s_{\delta}(x)$ only depend on the hypercube $A_{\delta}(x)$, so they take only finitely many values for $x\in[0,1]^{p}$. We now calculate the covariance
$\mathrm{Cov}(\tilde{m}^{(\varepsilon)}(x_{1}),\tilde{m}^{(\varepsilon)}(x_{2}))$ $=$ $\frac{1}{n}\mathbb{E}\Big{[}\mathbb{E}\left[\varepsilon^{2}|X\right]p_{\delta}(x_{1})^{-1}p_{\delta}(x_{2})^{-1}\mathbb{I}\{X\in A_{\delta}(x_{1})\}\mathbb{I}\{X\in A_{\delta}(x_{2})\}\Big{]}$
$=$ $\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}\frac{1}{n}\frac{s_{\delta}(x_{1})}{p_{\delta}(x_{1})^{2}}$
$=$ $\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}(n\tau_{\delta}(x_{1}))^{-1}.$
To obtain a process with unit variance, we consider
$(n\tau_{\delta}(x))^{1/2}\tilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\delta}$ (14)
for the empirical process $\mathbb{G}_{n}f=\sqrt{n}(P_{n}f-Pf)$ based on independent $(X_{1},\varepsilon_{1}),\dots,(X_{n},\varepsilon_{n})$ with distribution $P$ and empirical distribution $P_{n}$, indexed by $f\in\mathcal{F}_{\delta}$ with the finite function class
$\mathcal{F}_{\delta}:=\big{\{}f_{x,\delta}\colon[0,1]^{p}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\delta}(t,\epsilon)=\epsilon\,s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\},x\in[0,1]^{p}\big{\}}.$ (15)
We apply the result by Chernozhukov et al., 2014b formulated in Theorem A.1 in the Appendix A to approximate $\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\delta}$ with
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|.$
Therein, $B_{\delta}$ is a sequence of centered Gaussian processes indexed by $\mathcal{F}_{\delta}$ whose covariance function is given by
$\mathrm{Cov}(B_{\delta}(f_{x_{1},\delta}),B_{\delta}(f_{x_{2},\delta}))$ $=\mathbb{E}\left[f_{x_{1},\delta}(X_{1},\varepsilon_{1})f_{x_{2},\delta}(X_{1},\varepsilon_{1})\right]$
$=\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}.$ (16)
In particular, this implies $\mathrm{Var}(B_{\delta}(f))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables. The application of Theorem A.1 thus leads to the following theorem, which is proved in Section 4.
Theorem 3.1.
Let $\log\delta^{-1}=\mathcal{O}(\log n)$. Assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$. Then there exists a sequence of random variables
$\mathbf{S}_{\delta}\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|$
with independent standard normally distributed $(Z_{j})_{j\in\mathbb{N}}$ such that
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right).$
As mentioned above this direct approximation of the supremum avoids to approximate the whole empirical process uniformly by a Gaussian process. As a result we need weaker assumptions compared to more classical confidence bands derivations.
From the covariance calculation above and the assumptions on the covariate density and variance function it follows that the variance of the stochastic error is given by
$\mathrm{Var}\left(\tilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\delta}(x)}=\mathcal{O}\Big{(}\frac{1}{n\delta^{p}}\Big{)},$
i.e. the stochastic error profits from a larger $\delta$, which is caused by larger cells leading to more $\varepsilon_{j}$ being averaged and reducing the variance. Compared to the bound of the approximation error in (12), we recover the typical bias-variance trade-off in nonparametric statistics. The assumptions on $\delta$ in Theorem 2.1 lead to the commonly used undersmoothing. That means $\delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error allowing to build the confidence bands based on the distribution of the stochastic error.
4 Proofs
Subsequently we gather all the proofs starting with our main result Theorem 2.1. We will write $A\lesssim B$ equivalently for $A=\mathcal{O}(B)$. Moreover, we set $K:=\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $0,1]^{p}$.
4.1 Proof of Theorem 2.1 and its corollaries
Proof of Theorem 2.1.
We use the decomposition of the estimator and error from (11) and the definition of $\tilde{m}^{(\varepsilon)}$ in (13). For the sequence of random variables $\mathbf{S}_{\delta}$ as defined in Theorem 3.1 we obtain by the reverse triangle inequality that
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}$ $\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\mathbf{S}_{\delta}\bigg{|}\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
$T_{1}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\hat{m}^{(m)}(x)-m(x)\big{|},$ (17)
$T_{2}$ $=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|},$ (18)
$T_{3}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\tilde{m}^{(\varepsilon)}(x)-\hat{m}^{(\varepsilon)}(x)|.$ (19)
Subsequently, we will verify that all three terms are $o_{\mathbb{P}}((\log n)^{-1})$.
For $T_{1}$ from (17) note that the approximation error satisfies
$\hat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
$\qquad+m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1).$ (20)
Using (12), we get
$\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathfrak{d}(A_{\delta}(x))^{\alpha}\right]$
$=C_{H}\delta^{\alpha}p^{\alpha/2}.$ (21)
Since the number of cells is equal to $\delta^{-p}$, we obtain
$\mathbb{E}$ $\left[\sup_{x\in[0,1]^{p}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)|\right]$
$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\right]$
$\leq\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}\mathbb{E}\big{[}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\big{]}$
$=\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}(1-p_{\delta}(x))^{n}\leq\|m\|_{\infty}\delta^{-p}(1-c_{X}\delta^{p})^{n}.$
To handle (17), note that $n^{1/2}\sup_{x}|\tau_{\delta}(x)|^{1/2}\leq(C_{X}/c_{\sigma^{2}})^{1/2}n^{1/2}\delta^{p/2}$ by (8). Together with (20) and (21) we obain for all $\kappa>0$
$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\hat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/C_{X})^{1/2}}{n^{1/2}\delta^{p/2}\log n}\right)$
$\lesssim\mathbb{E}\big{[}\|m-\hat{m}^{(m)}\|_{\infty}\big{]}\kappa^{-1}(n\delta^{p})^{1/2}\log n$
$\lesssim(\delta^{2\alpha}n\delta^{p})^{1/2}\log n+\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n.$
The first term is $o(1)$ with (3) and for the second term we have
$\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n\lesssim\exp(-c_{X}n\delta^{p})\delta^{-p}(n\delta^{p})^{1/2}\log n\to 0$
because $\delta\leq n^{-1/(2\alpha+p)}$ by (3) and $\delta^{p}n\geq n^{2\nu}$ by (4). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.
For the term $T_{2}$ from (18) Theorem 3.1 yields
$T_{2}=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$ (22)
because $\nu\geq 4$. To see that the right-hand side is indeed $o_{\mathbb{P}}((\log n)^{-1})$, note that assumption (4) implies
$\frac{(\log n)^{5/2}}{\delta^{p/2}n^{1/2-1/\nu}}=\Big{(}\frac{(\log n)^{5}}{\delta^{p}n}n^{2/\nu}\Big{)}^{1/2}\to 0.$
The same argument applied to the other terms in (22) yields $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$.
Next we consider the remainder term $T_{3}$ from (19). By the definition of $\tilde{m}^{(\varepsilon)}$ in (13) and $\hat{p}_{\delta}(x)$ in (9) we obtain
$\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)=\tilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right).$
Together with (6) and $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$, we conclude
$T_{3}$ $\leq$ $\left(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in[0,1]^{p}}\left|\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right|$
$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-\mathbb{I}\{\hat{p}_{\delta}(x)=0\}\right|.$
To obtain the remainder rate $o_{\mathbb{P}}((\log n)^{-1})$, the second indicator can be written as $(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)$ and be handled as in (20). For the other term note that for fixed $x$ the random variable $B_{n}=n\hat{p}_{\delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\delta}(x)$, and we can apply Lemma 4.1. Together with (7), the union bound and Markov’s inequality we obtain
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\right|\mathbb{I}\{\hat{p}_{\delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
$\leq$ $\delta^{-p}\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}\delta^{p})^{q}\mathcal{O}((n\delta^{p})^{-3q/2})=\mathcal{O}\left(\delta^{-p}\frac{(\log n)^{3q/2}}{(\delta^{p}n)^{q/2}}\right)=o(1)$
which holds by assumption (4) for $q\in 2\mathbb{N}$, $q\geq\nu$. Hence, $T_{3}=o_{\mathbb{P}}((\log n)^{-1})$.
Altogether we have now proved
$\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|=\mathbf{S}_{\delta}+R_{n}$ (23)
for a remainder term $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$. Thus, the uniform coverage probability can be bounded from below by
$\mathbb{P}$ $\big{(}m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\big{)}$
$=\mathbb{P}\Big{(}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\leq c_{\delta}(\beta)\Big{)}$
$=\mathbb{P}\big{(}\mathbf{S}_{\delta}+R_{n}\leq c_{\delta}(\beta)\big{)}$
$\geq\mathbb{P}\big{(}\mathbf{S}_{\delta}\leq c_{\delta}(\beta)\big{)}-\mathbb{P}\big{(}c_{\delta}(\beta)-\kappa_{n}<\mathbf{S}_{\delta}\leq c_{\delta}(\beta)\big{)}-\mathbb{P}\left(R_{n}>\kappa_{n}\right),$
$\geq 1-\beta-\mathbb{P}\left(|\mathbf{S}_{\delta}-c_{\delta}(\beta)|\leq\kappa_{n}\right)-\mathbb{P}\left(R_{n}>\kappa_{n}\right),$
where the last term converges to zero for $\kappa_{n}=(\log n)^{-1}$ and the probability in the middle term can be upper bounded by
$\sup_{\xi\in\mathbb{R}}\mathbb{P}\left(|\mathbf{S}_{\delta}-\xi|\leq\kappa_{n}\right)$ $\leq 4\kappa_{n}\left(\mathbb{E}\left[\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|\right]+1\right)$
$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\right]+1\right)$
$=\kappa_{n}\mathcal{O}(\sqrt{\log n})=o(1)$
which follows from Corollary 2.1 by Chernozhukov et al., 2014a and (6). Thus we obtain
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.\qed$
Proof of Corollary 2.2.
To verify the uniform rate, we first note that assumption (3) was only required to ensure that the approximation error $T_{1}$ from (17) is of order $o_{\mathbb{P}}((\log n)^{-1})$. The estimates following (20) and (8) show that the approximation error satisfies
$\mathbb{E}\big{[}\sup_{x\in[0,1]^{p}}|\hat{m}^{(m)}(x)-m(x)|\big{]}\lesssim n^{-1/2}\delta^{-p/2}\mathbb{E}[T_{1}]\lesssim\delta^{\alpha}+\delta^{-p}(1-c_{X}\delta^{p})^{n},$
where the second term is negligible if $\log(\delta^{-1})=o(n\delta^{p})$. The latter condition is implied by assumption (4). Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (6) and (8), we conclude $\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}})$. Choosing $\delta=(n/\log n)^{-1/(2\alpha+p)}$ balances both terms and yields the optimal rate of convergence. ∎
Proof of Corollary 2.3.
As in the proof of Theorem 2.1 we need to show $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$ for the additional remainder term
$R_{n}$ $=$ $\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in[0,1]^{p}}\hat{\tau}_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\bigg{|}$
$\leq$ $\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)^{1/2}-\tau_{\delta}(x)^{1/2}|}{\tau_{\delta}(x)^{1/2}}$
$=$ $(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1}))\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)-\tau_{\delta}(x)|}{\tau_{\delta}(x)^{1/2}(\tau_{\delta}(x)^{1/2}+\hat{\tau}_{\delta}(x)^{1/2})}$
$\leq$ $\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}((\log n)^{-1})$
by (23), (25), (6) and the assumption on $\hat{\tau}_{\delta}$. ∎
4.2 Proof of Theorem 3.1
Proof of Theorem 3.1.
We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem A.1 in the appendix to fix the notation. While the function class in this result does not depend on $n$ and $\delta$, the theorem can be applied to $\mathcal{F}_{\delta}$ for every $n$ or $\delta$, respectively, due to the universality of the occuring constants. As a consequence we get a sequence $\mathbf{S}_{\delta}$ of approximating random variables as pointed out by Chernozhukov et al., 2014b (, Remark 2.1). We consider the function class $\mathcal{F}_{\delta}$ from (15). To obtain the claimed result for the absolute value, we need to consider $\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}$. Due to Chernozhukov et al., 2014b (, Corollary 2.1) we consider $\mathcal{F}_{\delta}$ without loss of generality.
The functions $f_{x,\delta}$ are measurable and thus the function class is pointwise measurable. To see that $\mathcal{F}_{\delta}$ is a VC-type class as defined in the appendix, we calculate a measurable envelope $F$:
$\sup_{x\in[0,1]^{p}}|f_{x,\delta}(t,\epsilon)|$ $=\sup_{x\in[0,1]^{p}}|\epsilon|s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\}$
$\leq|\epsilon|(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}=:F(t,\epsilon)$ (24)
applying $c_{\sigma^{2}}c_{X}\delta^{p}\leq s_{\delta}(x)\leq C_{\sigma^{2}}C_{X}\delta^{p}$ from (10), (1) and (2). The finite size $|\mathcal{F}_{\delta}|=\delta^{-p}$ is an upper bound for any covering number of $\mathcal{F}_{\delta}$. Hence,
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\delta^{-p}\leq\delta^{-p}/\kappa$
for all $\kappa\in(0,1]$, $\mathcal{F}_{\delta}$ is indeed of VC-type and we can choose $A=\delta^{-p}$ and $v=1$ in Theorem A.1.
We proceed by verifying the conditions on the moments. Since
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{2}=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}\right]=1,$
$\tilde{\sigma}$ from Theorem A.1 is equal to $1$. Moreover,
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{3}$ $=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}|^{3}\right]$
$\leq K(c_{\sigma^{2}}c_{X})^{-3/2}C_{X}\delta^{-p/2}.$
For the envelope $F$ from (24) we obtain
$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
We thus choose $b$ from Theorem A.1 as
$b:=\delta^{-p/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}})^{-3/2}C_{X},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}})^{-1/2}\right\}.$
Since $b$ is of order $\delta^{-p/2}$, we have $b\geq 1$ for $\delta$ small enough. To apply Theorem A.1 we obtain
$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\tilde{\sigma}))=c(\log n\vee\log(\delta^{-p}\delta^{-p/2}C_{b}))$
$\lesssim\log n\vee p\log(\delta^{-1})\lesssim\log n.$
Let $B_{\delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem A.1 yields that there exists a random variable
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}}B_{\delta}(f)=\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|=\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|$
such that
$\mathbb{P}$ $\Big{(}\Big{|}\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|-\mathbf{S}_{\delta}\Big{|}>\frac{b\Gamma_{n}(\log n)^{1/2}}{n^{1/2-1/\nu}}+\frac{(b\log n)^{1/2}\Gamma_{n}^{3/4}}{n^{1/4}}+\frac{(b\Gamma_{n}^{2}\log n)^{1/3}}{n^{1/6}}\Big{)}$
$\lesssim(\log n)^{-1}+\frac{\log n}{n}.$
In view of (14), we conclude
$\bigg{|}$ $\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}$
$=\Big{|}\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|-\mathbf{S}_{\delta}\Big{|}$
$=\mathcal{O}_{\mathbb{P}}\left(\frac{b\Gamma_{n}(\log n)^{1/2}}{n^{1/2-1/\nu}}+\frac{(b\log n)^{1/2}\Gamma_{n}^{3/4}}{n^{1/4}}+\frac{(b\Gamma_{n}^{2}\log n)^{1/3}}{n^{1/6}}\right)$
$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$
since $\Gamma_{n}=\mathcal{O}(\log n)$ and $b=\mathcal{O}(\delta^{-p/2})$. Finally note, that the structure of the covariance (16) implies that
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|,$ (25)
for independent standard normally distributed $Z_{1},Z_{2},\dots$. ∎
4.3 Moment bounds for binomial distributions
We need a uniform bound on centered moments of binomial random variables. To this end, we use Skorski, (2025, Theorem 4), which gives for any random variable $B\sim\text{Bin}(n,p)$ with $n\in\mathbb{N}$, $p\in(0,1)$ and for any $q>1$ that
$\mathbb{E}\big{[}(B-np)^{q}\big{]}^{1/q}=C(n,p,q)\max\big{\{}k^{1-k/q}(np(1-p))^{k/q}:k=1,\ldots,\lfloor q/2\rfloor\big{\}},$
where $C(n,p,q)$ is uniformly bounded by $(3e)^{-1}\leq C(n,p,q)\leq(5/2)^{1/5}e^{1/2}$. If $np>1$, we obtain
$\max\{k^{1-k/q}(n\sigma^{2})^{k/q}:k=1,\ldots,\lfloor q/2\rfloor\}\leq(np)^{\lfloor q/2\rfloor/q}\max\{k^{1-k/q}:k=1,\ldots,\lfloor q/2\rfloor\}.$
Hence, for fixed $q>1$ we have
$\mathbb{E}[(B-np)^{q}]\lesssim(np)^{q/2}\text{ uniformly in }n\text{ and }p.$ (26)
Lemma 4.1.
Let $B\sim\text{Bin}(n,p)$ such that $p^{-1}\lesssim n$. For any fixed integer $q>1$ it holds that
$\mathbb{E}\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$ $\lesssim(np)^{-3q/2}.$
Proof.
Using simple calculations one gets
$\frac{1}{B}-\frac{1}{np}=\frac{np-B}{(np)^{2}}+\frac{(np-B)^{2}}{(np)^{3}}+\frac{(np-B)^{3}}{B(np)^{3}}.$ (27)
Hence we get
$\mathbb{E}$ $\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$
$\lesssim\mathbb{E}\left[\left(\frac{np-B}{(np)^{2}}\right)^{q}\right]+\mathbb{E}\left[\left(\frac{(np-B)^{2}}{(np)^{3}}\right)^{q}\right]+\mathbb{E}\left[\mathbb{I}\{B>0\}\left(\frac{(np-B)^{3}}{B(np)^{3}}\right)^{q}\right]$
$\leq(np)^{-2q}\mathbb{E}\left[\left(np-B\right)^{q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{2q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{3q}\right].$
We apply (26) in the case where $p^{-1}=\mathcal{O}(n)$. We obtain
$\mathbb{E}$ $\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$
$\lesssim(np)^{-2q}\mathbb{E}\left[(np-B)^{q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{2q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{3q}\right]$
$\lesssim(np)^{-3q/2}+(np)^{-2q}+(np)^{-3q/2}\lesssim(np)^{-3q/2},$
which yields the claim. ∎
Lemma 4.2.
If (4) is satisfied, we have
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}\big{(}(\log n)^{-3/2})\big{)}.$
Proof.
Note that
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}$ $=$ $\sup_{x\in[0,1]^{p}}\frac{1}{np_{\delta}(x)}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}.$
Since $x\mapsto\hat{p}_{\delta}(x)$ is piecewise constant with at most $\delta{-p}$ different values, the supremum can be handled with a union bound. Moreover, we can exploit that $\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}$ is $Bin(n,p_{\delta}(x))$-distributed. The binomial moment result (26) and (7) yield for any $q\in\mathbb{N}$ that
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
$\qquad\leq\delta^{-p}\frac{1}{(nc_{X}\delta^{p})^{q}}\mathbb{E}\Big{[}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}^{q}\Big{]}\frac{(\log n)^{3q/2}}{\kappa^{q}}$
$\qquad\lesssim\frac{\delta^{-p}(\log n)^{3q/2}}{\kappa^{q}(\delta^{p}n)^{q/2}}=o(1)$
which holds by assumption (4) for large enough $q$. ∎
Appendix A Approximation of suprema of empirical processes
Denote by $P_{n}$ the empirical distribution of independent random variables $Z_{1},\dots,Z_{n}$ with distribution $P$. Then we use the empirical process notation $\mathbb{G}_{n}f=n^{1/2}(P_{n}f-Pf)$ indexed in $f\in\mathcal{F}$, where we assume that the function class $\mathcal{F}$ is of VC-type with envelope $F$. This means that there exist constants $A,v>0$ such that for the covering number
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F},\|\cdot\|_{Q,2},\varepsilon\|F\|_{Q,2})\leq(A/\varepsilon)^{v}\qquad\text{for all}\;\varepsilon\in(0,1]$
is valid. Here we use the notation $\|f\|_{Q,2}=(Qf^{2})^{1/2}$ and $\mathcal{Q}$ denotes the set of all finitely discrete probability measures. The next theorem is Corollary 2.2 by Chernozhukov et al., 2014b .
Theorem A.1.
Suppose that $\mathcal{F}$ is a pointwise measurable function class with envelope $F$ that is VC-type for constants $A\geq e,v\geq 1$. Let $P$ denote the distribution of $Z_{1}$ and let $G_{P}$ be a centered Gaussian process indexed in $\mathcal{F}$ with
$\mathbb{E}[G_{P}(f)G_{P}(g)]=P(fg)=\mathbb{E}[f(Z_{1})g(Z_{1})].$
Suppose also that for some $b\geq\tilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\tilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$. Then for every $\gamma\in(0,1)$ there exists a random variable $\mathbf{S}\overset{d}{=}\sup_{f\in\mathcal{F}}G_{P}f$ such that
$\mathbb{P}\Big{(}\Big{|}\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big{|}>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\tilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\tilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big{)}\leq C\Big{(}\gamma+\frac{\log n}{n}\Big{)}$
where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\tilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).

### v2

@@
-Asymptotic confidence bands for the histogram regression estimator
+arXiv:2508.12391v2[math.ST] 27 Feb 2026
 Abstract
-In a multivariate nonparametric regression model with a Hölder continuous regression function and heteroscedastic noise asymptotic uniform confidence bands are constructed based on the histogram estimator.
-The radius of the confidence bands does not depend on an extreme value distribution, but instead can be numerically calculated for the chosen binning.
-Keywords: Multivariate nonparametric regression, histogram, confidence bands, heteroscedastic noise
+Asymptotic uniform confidence bands are constructed for a multivariate nonparametric regression model with heteroscedastic noise, employing histogram estimators under flexible partition conditions.
+The construction is especially applicable to unsmooth regression functions of Hölder regularity less than one.
+While the radius of the confidence bands could be approximated via the Gumbel distribution, our construction does not depend on an extreme value distribution, but instead can be explicitly calculated for the chosen partition.
+Keywords: Multivariate nonparametric regression, partitioning estimator, confidence bands, heteroscedastic noise
 MSC 2020: 62G08, 62G20, 62H12
 1 Introduction
-We consider classical histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise, and derive asymptotic confidence bands.
+We consider histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise.
+Our aim is to derive asymptotic histogram based confidence bands.
 There is a substantial amount of literature on confidence bands based on different nonparametric regression estimators.
@@
 There are also bootstrap confidence bands for nonparametric regression, see for instance Hall, (1993), Neumann and Polzehl, (1998), Claeskens and Van Keilegom, (2003) or Hall and Horowitz, (2013).
+Based on spline estimators regression confidence bands were constructed by Wang and Yang, (2009) in the random covariate case, and by Krivobokova et al., (2010) in the deterministic design case.
+Cai et al., (2019) considered confidence bands for the mean and the variance function for fixed design.
+These confidence bands are based on Nadaraya-Watson-type estimators, where for the variance case the residuals are estimated based on spline-estimators.
 Sabbah, (2014) has shown confidence bands for quantile estimators which is an alternative to a regression model, and Birke et al., (2010) have proved confidence bands in an inverse regression model.
 The aforementioned articles only considered univariate regression.
+For multidimensional covariates often semiparametric regression models are considered due to the curse of dimensionality.
+For instance, confidence bands for the link function in a single-index model were developed by Gu and Yang, (2015), for additive regression functions by Härdle et al., (2015), and for additive semi-parametric regression functions by Cattaneo et al., (2024).
+In contrast we focus on multivariate nonparametric regression models without additional structural assumptions.
 In the multivariate case Konakov and Piterbarg, (1984) have analyzed the asymptotic distribution of the maximal deviation for the Nadaraya-Watson estimate in a random design setting.
@@
 Chao et al., (2017) have covered confidence bands for multivariate quantile and expectile regression with mean regression as a special case.
-Despite the popularity and widespread application of histogram estimators, the only uniform confidence band result for a histogram estimator that we are aware of is by Smirnov, (1950) and covers univariate density estimation.
-We want to close this gap by presenting a confidence band result in a multivariate regression model.
-In an undersmooting regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$.
-The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,J}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $J$ is given by the number of cells in the partition of the histogram.
-In particular, our confidence bands do not rely on an extreme value distribution.
+Due to their easy and efficient implementation and clear interpretability histogram regression estimators (also called regressogram) enjoy a high popularity and widespread application, see e.g.
+Györfi et al., (2002); Wasserman, (2006).
+Also in the development of advanced statistical methodologies, histogram estimators continue to serve as a fundamental building block, e.g.
+for regression trees Breiman et al., (1984) or in the context of privacy, see Berrett et al., (2021).
+In a density estimation setting the study of confidence bands goes back to Smirnov, (1950).
+The regression histogram estimators are the simplest case of partitioning-based estimators whose asymptotic properties are studied in depth by Cattaneo et al., (2020).
+The latter article also includes confidence bands with a focus on general (bias corrected) partitioning-based estimators.
+For the case of multivariate histogram estimators the applicability of their construction is however limited.
+We want to close this gap by presenting an easy to implement confidence band result in a multivariate regression model for histogram estimators under very mild conditions on the partition.
+This especially allows us to incorporate a-priori knowledge on the regression function into the construction of the partition to improve the finite sample performance.
+In an undersmoothing regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$.
+Note that all above mentioned articles on confidence bands require the regression function to be at least continuously differentiable while we will allow for regression functions which are less than Lipschitz regular.
+The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,\Delta}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $\Delta$ is given by the number of cells of the partition of the histogram.
+Whereas our confidence intervals do not need to be based on an extreme value distribution, our critical values can be related to the quantiles of the Gumbel distribution.
 We also incorporate the estimation of the unknown variance and a possible unknown distribution of the covariates.
 We obtain a confidence band result for the simple case of a Hölder continuous regression function for every covariate dimension, without a continuity assumption on the covariate density.
-In contrast for smooth nonparametric regression estimators typically it is assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density have to increase, see Chao et al., (2017), for instance.
-Instead we apply error moment existence, but the assumptions are not restrictive in the sense of exponential moments.
+In contrast, for smooth nonparametric regression estimators it is typically assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density has to increase, see, for instance, Chao et al., (2017).
+Instead we only exploit finite higher moments of the error distribution, an assumption which is especially satisfied for finite exponential moments.
 A central part of the proof for the histogram based confidence band is the application of a result by Chernozhukov et al., 2014b regarding the Gaussian approximation for the suprema of empirical processes.
@@
 This method has also been applied by Patschkowski and Rohde, (2019) to construct adaptive confidence bands for probability densities.
-It is an important difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
-The main result is presented in Section 2. The proof strategy is outlined in Section 3, while detailed proofs are postponed to Section 4.
+It is a key difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
+The main result is presented and discussed in Section 2.
+The proof strategy is outlined in Section 3.
+In Section 4 we show a numerical illustration.
+Some conclusions are given in Section 5.
+Detailed proofs are postponed to Section 6 and in Appendix A extensions are discussed.
 2 Histogram estimator and confidence bands
-On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in[0,1]^{p}$ and $Y\in\mathbb{R}$.
-The goal is to estimate the regression function $m\colon[0,1]^{p}\to\mathbb{R}$ with
+On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in\mathcal{X}$, $Y\in\mathbb{R}$ and where $\mathcal{X}\subseteq\mathbb{R}^{p}$ is some bounded feature space.
+The goal is to estimate the regression function $m\colon\mathcal{X}\to\mathbb{R}$ with
 $Y=m(X)+\varepsilon$
@@
 We further assume that the density $f_{X}$ of the covariate or feature vector $X$ and the variance function are bounded from below and above, i.e.
-$0<c_{X}\leq f_{X}(\cdot)\leq C_{X}$ (1)
+$0<c_{X}\leq f_{X}(\cdot)\leq C_{X},$ (1)
 $0<c_{\sigma^{2}}\leq\sigma^{2}(\cdot)\leq C_{\sigma^{2}}$ (2)
@@
 The histogram estimator relies on a local constant approximation of $m$.
-For a given partition of the feature space $0,1]^{p}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition.
-We will choose an equidistant grid to construct the partition such that all cells are hypercubes of the same size.
-For $\delta^{-1}\in\mathbb{N}$ we consider the grid with edge length $\delta$ on $0,1]^{p}$.
-Here $\delta=\delta_{n}$ is a sequence depending on the sample size with $\delta\to 0$ for $n\to\infty$.
-For $x\in[0,1]^{p}$ let $A_{\delta}(x)$ be the hypercube with volume $\delta^{p}$ that contains $x$.
-Note that the equidistant grid is considered for simplicity and all results carry over to more irregular partitions as long as the volume and the diameters of all cells in the same order.
-The histogram estimator is then defined by
-$\hat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
-where $\mathbb{I}\{\cdot\}$ denotes the indicator function. If there is no observation in $A_{\delta}(x)$ we define the estimator as zero on this cell.
+For a given partition of the feature space $\mathcal{X}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition.
+We will impose the following regularity assumption on the partition:
+Assumption 2.1.
+Let $\bigcup_{k=1}^{\Delta}A_{k}=\mathcal{X}$ be a partition of $\mathcal{X}$ of size $\Delta=\Delta_{n}\in\mathbb{N}$ into disjoint subsets $A_{k}\subset\mathcal{X},k=1,\dots,\Delta$, satisfying for some constants $c_{p},C_{p}>0$ that
+$c_{p}\Delta^{-1}\leq\operatorname{vol}(A_{k})\leq C_{p}\Delta^{-1}\text{ and }\operatorname{diam}(A_{k})\leq C_{p}\Delta^{-1/p}\text{ for all }k=1,\dots,\Delta$ (3)
+with volume $\operatorname{vol}(A_{k})$ and diameter $\operatorname{diam}(A_{k}):=\sup_{x_{1},x_{2}\in A_{k}}\|x_{1}-x_{2}\|$ of $A_{k}$.
+On the one hand, the assumption guarantees that the overall volume $\operatorname{vol}(\mathcal{X})$ is roughly equally distributed over the cells of the partition.
+The condition on diameter ensures that the cells of the partition do not degenerate in any direction, but instead spread similarly in all directions.
+On the other hand, Assumption 2.1 allows for a flexible choice of the partition such that, e.g.
+prior knowledge of the regression function could be incorporated in the construction to improve the finite sample performance.
+In Section 4 we will see a numerical example in this direction.
+For $x\in\mathcal{X}$ let $A_{\Delta}(x)$ be the subset $A_{k}$ that contains $x$. The histogram estimator is then defined by
+$\widehat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}$ (4)
+where $\mathbb{I}\{\cdot\}$ denotes the indicator function.
+If there is no observation in $A_{\Delta}(x)$ we define the estimator as zero on this cell.
+The simplicity of histogram estimators comes with some computational advantages.
+First, they have lower computational costs than kernel estimators when many evaluations are needed ($O((m\wedge\Delta)n)$ instead of $O(mn)$ for $m$ point evaluations).
+Second, they serve as summary statistics to compress data and reduce storage space from $n$ observations to only $\Delta$ cell-wise means.
 In the following theorem we present a confidence band result based on the histogram estimator.
-Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $0,1]^{p}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
+Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $m\colon\mathcal{X}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
 Theorem 2.2.
-Let $\alpha\in(0,1],C_{H}>0$ and assume that
-$n\delta^{2\alpha+p}(\log n)^{2}\to 0.$ (3)
-Further let $4\leq\nu\in\mathbb{N}$ be such that $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ and
-$\frac{(\log n)^{5}}{\delta^{p}n^{1-2/\nu}}$ $\to 0$ (4)
-and assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$.
+Let the partition satisfy Assumption 2.1. Let $\alpha\in(0,1],C_{H}>0$ and assume that
+$n\Delta^{-1-2\alpha/p}(\log n)^{2}\to 0.$ (5)
+Assume that $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that
+$\frac{\Delta(\log n)^{5}}{n^{1-2/\nu}}$ $\to 0.$ (6)
 For independent standard normally distributed random variables $(Z_{j})_{j\in\mathbb{N}}$ and some $\beta\in(0,1)$ let $c_{\Delta}(\beta)>0$ be given by
-$\mathbb{P}\left(\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\leq c_{\delta}(\beta)\right)=1-\beta.$
+$\mathbb{P}\left(\max_{j=1,\ldots,\Delta}|Z_{j}|\leq c_{\Delta}(\beta)\right)=1-\beta.$ (7)
 For
@@
 the confidence band
-$\mathcal{C}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2}\right]$
+$\mathcal{C}_{n}(x)=\left[\widehat{m}(x)-c_{\Delta}(\beta)(\tau_{\Delta}(x)n)^{-1/2},\widehat{m}(x)+c_{\Delta}(\beta)(\tau_{\Delta}(x)n)^{-1/2}\right]$
 satisfies
-$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
+$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.$
 The proof strategy is explained in Section 3, and the proofs are given in Section 6.
@@
 Note that by Gaussian concentration, we have
-$\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2p\log\delta^{-1}})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (6)
-due to $\log\delta^{-1}\lesssim\log n$ which is implied by assumption (4). Therefore, $c_{\delta}(\beta)\lesssim\sqrt{\log n}$.
+$\max_{j=1,\ldots,\Delta}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2\log\Delta})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (9)
+due to $\log\Delta\lesssim\log n$ which is implied by assumption (6).
+Therefore, $c_{\Delta}(\beta)\lesssim\sqrt{\log n}$.
+Moreover, the quantile $c_{\Delta}(\beta)$ can be explicitly represented in terms of the quantile function of the standard normal distribution since
+$1-\beta=\mathbb{P}\left(\max_{j=1,\ldots,\Delta}|Z_{j}|\leq c_{\Delta}(\beta)\right)=\mathbb{P}\left(|Z_{1}|\leq c_{\Delta}(\beta)\right)^{\Delta}=(2\Phi(c_{\Delta}(\beta))-1)^{\Delta}.$
+with cdf $\Phi$ of $\mathcal{N}(0,1)$. Consequently, $c_{\Delta}(\beta)=\Phi^{-1}(\frac{1}{2}(1+(1-\beta)^{1/\Delta}))$ is easy to compute.
 Define
 $p_{\Delta}(x)=\mathbb{P}(X\in A_{\Delta}(x))=\int_{A_{\Delta}(x)}f_{X}(t)\,dt.$
-By (1) and (2) we have
-$c_{X}\delta^{p}\leq p_{\delta}(x)\leq C_{X}\delta^{p}$ (7)
+By (1) and (2) we obtain
+$c_{X}\operatorname{vol}(A_{\Delta}(x))\leq p_{\Delta}(x)\leq C_{X}\operatorname{vol}(A_{\Delta}(x))$ (10)
 and
-$\frac{c_{X}}{C_{\sigma^{2}}}\delta^{p}\leq\tau_{\delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\delta^{p}.$ (8)
-Therefore, the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)/(n\delta^{p})}$ uniformly in $x\in[0,1]^{p}$.
+$\frac{c_{X}}{C_{\sigma^{2}}}\operatorname{vol}(A_{\Delta}(x))\leq\tau_{\Delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\operatorname{vol}(A_{\Delta}(x)).$ (11)
+Due to Assumption 2.1 we have $\operatorname{vol}(A_{\Delta}(x))=\mathcal{O}(\Delta^{-1})$ uniformly in $x\in\mathcal{X}$ and thus the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)\Delta/n}$ uniformly on $\mathcal{X}$.
+Example 2.3.
+Let $\mathcal{X}=[0,1]^{p}$ and consider a partition based on an equidistant grid such that all cells $(A_{k})_{k=1,\dots,\Delta}$ are hypercubes of the same size.
+Denoting the edge length of the cubes by $\delta>0$, we have $\Delta=\delta^{-p}$, $\operatorname{diam}(A_{k})=\sqrt{p}\delta=\sqrt{p}\Delta^{-1/p}$ and $\operatorname{vol}(A_{k})=\delta^{p}=\Delta^{-1}$.
 In the homoscedastic case with constant $\sigma^{2}(\cdot)\equiv\sigma^{2}$ we obtain $\tau_{\Delta}(x)=p_{\Delta}(x)/\sigma^{2}$.
-For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\delta}(x)=\delta^{p}$ and in that case the confidence band simplifies to
-$\mathcal{C}_{n}(x)=\Big{[}\hat{m}(x)-c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}},\hat{m}(x)+c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}}\Big{]}.$
+For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\Delta}(x)=\Delta^{-1}$ and in that case the confidence band simplifies to
+$\mathcal{C}_{n}(x)=\Big[\widehat{m}(x)-c_{\Delta}(\beta)\sigma\sqrt{\frac{\Delta}{n}},\widehat{m}(x)+c_{\Delta}(\beta)\sigma\sqrt{\frac{\Delta}{n}}\Big].$
 The assumptions (5) and (6) in Theorem 2.2 are up to logarithms that
-$n\delta^{p+2\alpha}\to 0\text{ and }n\delta^{p}/(n^{2/\nu})\to\infty,$
+$n/\Delta^{1+2\alpha/p}\to 0\text{ and }n/(\Delta n^{2/\nu})\to\infty,$
 where $2/\nu$ is arbitrary small if all moments of $\varepsilon$ exist.
-Since the approximation error of the histogram estimator is of order $\delta^{\alpha}$ while the stochastic error term is of order $(n\delta^{p})^{-1/2}$, the first assumption is an undersmoothing condition.
+Since the approximation error of the histogram estimator is of order $\Delta^{-\alpha/p}$ while the stochastic error term is of order $(\Delta/n)^{1/2}$, the first assumption is an undersmoothing condition.
 The second condition implies that the average number of observations in all cells goes to infinity such that the estimator is consistent.
@@
 However, the assumptions are not contradictory, and for every dimension $p$ no bias reduction is needed.
-In contrast, the confidence bands by Chao et al., (2017) require increasing smoothness orders for the regression function $m$ and covariate density $f_{X}$ if the dimension $p$ increases.
-We are not aware of any other confidence band result in the literature for the histogram estimator in the multivariate or univariate regression case.
-However, the structure of the histogram with fixed deterministic cells might imply that the difference in the proof from the univariate to the multivariate case is minor.
-The first result for the histogram by Smirnov, (1950) studies univariate density estimation.
-In addition to the different estimation problem, this result differs from our result in another crucial aspect.
-The asymptotic bands rely on an extreme value distribution that is independent of $\delta$.
-Instead our result provides bands with explicit dependence on $\delta$.
-Thus, the knowledge of the limit distribution is not necessary, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution.
+The special case of confidence bands based on a histogram with a cubic partition is also covered by Cattaneo et al., (2020, Theorem 6.1 in combination with Remark 6.2).
+For general partitions in dimensions $p\geq 2$ Theorem 6.2 in Cattaneo et al., (2020) requires more regular regression functions compared to Theorem 2.2.
+More precisely, they need $\Delta^{3}/n\to 0$ (in our notation) instead of (6).
+An optimal choice $\Delta=n^{p/(2\alpha+p)}$ (up to logarithms, see Corollary 2.5) satisfies $\Delta^{3}/n\to 0$ only for $\alpha>p$.
+Cattaneo et al., (2020) also assume differentiability of the regression function.
+Theorem 2.2 does not require that the regression function’s and covariate density’s regularities have to grow with the dimension, which is a common restriction in the literature on nonparametric regression with multivariate covariates, see Proksch, (2016) and Chao et al., (2017) for confidence bands, and Dette and Derbort, (2001) and Neumeyer and Van Keilegom, (2010), among others, for different estimation and testing problems.
+Typically, confidence bands regression functions rely on an extreme value distribution that is independent of $\Delta$.
+Based on Theorem 2.2 and the Gumbel extreme value distribution, we can derive an analogous result.
+Defining
+$a_{\Delta}$ $=$ $\sqrt{2\log\Delta}\qquad\text{and}$
+$b_{\Delta}$ $=$ $\sqrt{2\log\Delta}-\frac{\log(\log\Delta)+\log(4\pi)-\log(4)}{\sqrt{8\log\Delta}},$
+the critical value $c_{\Delta}(\beta)$ can be approximated by
+$\widetilde{c}_{\Delta}(\beta)=b_{\Delta}-\frac{\log(-\log(1-\beta))}{a_{\Delta}}.$ (12)
+Let the confidence interval $\widetilde{\mathcal{C}}_{n}(x)$ be defined as $\mathcal{C}_{n}(x)$, but replacing $c_{\Delta}(\beta)$ by $\widetilde{c}_{\Delta}(\beta)$.
+The proof of the corollary is given in Section 6.
+Corollary 2.4.
+Under the assumptions of Theorem 2.2 it holds that
+$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\widetilde{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\right)=1-\beta.$
+In contrast to the corollary Theorem 2.2 provides bands with explicit dependence on $\Delta$, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution.
+Indeed, the approximation accuracy of $\widetilde{c}_{\Delta}(\beta)$ increases only slowly in the number of cells $\Delta$ and consistently overestimates $c_{\Delta}(\beta)$ as shown in Figure 1.
 Applying an approximation by the maximum of Gaussian random variables instead of a limit distribution is also suggested by Kreiß et al., (2026) for confidence bands for the spectral density in time series analysis.
-From Theorem 2.1 we obtain a uniform rate of convergence of the histogram estimator.
+[…]
+Figure 1: Absolute error $\widetilde{c}_{\Delta}(\beta)-c_{\Delta}(\beta)$ and relative error $(\widetilde{c}_{\Delta}(\beta)-c_{\Delta}(\beta))/c_{\Delta}(\beta)$ for $\beta=0.05$ in dependence of $\Delta$.
+From Theorem 2.2 we obtain a uniform rate of convergence of the histogram estimator. Then we obtain the following corollary whose proof is given in Section 6.
 Corollary 2.5.
-Assume $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}/(\delta^{p}n^{1-2/\nu})\to 0$.
+Let the partition satisfy Assumption 2.1.
+Assume $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}\Delta/n^{1-2/\nu}\to 0$.
 Then we have
-$\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}}\right).$
-For $\nu>2+p/\alpha$ and $\delta=(n/\log n)^{-1/(2\alpha+p)}$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
+$\sup_{x\in\mathcal{X}}|\widehat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\Delta^{-\alpha/p}+\sqrt{(\log n)\Delta/n}\right).$
+For $\nu>2+p/\alpha$ and $\Delta=\lfloor(n/\log n)^{p/(2\alpha+p)}\rfloor$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
 This convergence rate coincides the minimax-optimal rate for the uniform loss, cf. Tsybakov, (2009). The proof is given in Section 6.
 We will now consider the case where the covariate distribution and the variance function of the noise are unknown.
-To this end, let $\hat{\tau}_{\delta}(x)$ be an estimator for $\tau_{\delta}(x)$ from (5) and define the confidence band as
-$\hat{\mathcal{C}}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2}\right].$
+To this end, let $\widehat{\tau}_{\Delta}(x)$ be an estimator for $\tau_{\Delta}(x)$ from (8) and define the confidence band as
+$\widehat{\mathcal{C}}_{n}(x)=\left[\widehat{m}(x)-c_{\Delta}(\beta)(\widehat{\tau}_{\Delta}(x)n)^{-1/2},\widehat{m}(x)+c_{\Delta}(\beta)(\widehat{\tau}_{\Delta}(x)n)^{-1/2}\right].$
 Corollary 2.6.
-Under the assumptions of Theorem 2.1 and
-$\sup_{x\in[0,1]^{p}}\left|\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2})$
+Under the assumptions of Theorem 2.2 and if
+$\sup_{x\in\mathcal{X}}\left|\frac{\widehat{\tau}_{\Delta}(x)}{\tau_{\Delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2}),$
 we have
-$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\hat{\mathcal{C}}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
+$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\widehat{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.$
 The proof is again postponed to Section 6.
-To construct the estimator $\hat{\tau}_{\delta}(x)$, we first consider the homoscedastic case where $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$ can be estimated by $\hat{\tau}_{\delta}(x)=\hat{p}_{\delta}(x)/\hat{\sigma}^{2}$ with
-$\hat{p}_{\delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}.$ (9)
-Note that $\hat{p}_{\delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution.
+To construct the estimator $\widehat{\tau}_{\Delta}(x)$, we first consider the homoscedastic case where $\tau_{\Delta}(x)=p_{\Delta}(x)/\sigma^{2}$ can be estimated by $\widehat{\tau}_{\Delta}(x)=\widehat{p}_{\Delta}(x)/\widehat{\sigma}^{2}$ with
+$\widehat{p}_{\Delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}.$ (13)
+Note that $\widehat{p}_{\Delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution.
 This estimator has the desired rate which we show in Lemma 6.2.
 The constant variance can be estimated using residuals based on the histogram regression estimator, i.e.
-$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-\hat{m}(X_{i})\big{)}^{2}.$
+$\widehat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big(Y_{i}-\widehat{m}(X_{i})\big)^{2}.$
 Other variance estimators under homoscedasticity and variance function estimators under heteroscedasticity are considered by Shen et al., (2020), among others.
 Nevertheless, one can also apply a histogram estimator for the variance function $\sigma^{2}(x)=\mathbb{E}[Y^{2}|X=x]-m(x)^{2}$ defined by
-$\hat{\sigma}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}-\hat{m}(x)^{2}$
-$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\hat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
-Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(((\log n)/(n\delta^{p}))^{1/2})$ holds as in Corollary 2.2 under the same assumptions for observations $(X,Y^{2})$, in particular moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$.
-The histogram estimator $\hat{\sigma}^{2}(\cdot)$ is constant on each hypercube $A_{\delta}(x)$ and thus
+$\widehat{\sigma}_{\Delta}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}-\widehat{m}(x)^{2}$
+$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\widehat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}.$
+Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(\sqrt{(\log n)\Delta/n})$ holds as in Corollary 2.5 under the same assumptions for observations $(X,Y^{2})$, in particular under moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$.
+The histogram estimator $\widehat{\sigma}_{\Delta}^{2}(\cdot)$ is constant on each subset $A_{\Delta}(x)$ and thus
 $s_{\Delta}(x)=\mathbb{E}\big[\sigma^{2}(X)\mathbb{I}_{A_{\Delta}(x)}(X)\big]=\int_{A_{\Delta}(x)}\sigma^{2}(t)f_{X}(t)\,dt$ (14)
-is approximated by $\int_{A_{\delta}(x)}\hat{\sigma}^{2}(x)f_{X}(t)\,dt=\hat{\sigma}^{2}(x)p_{\delta}(x)$. Hence, we can estimate $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$ by
-$\hat{\tau}_{\delta}(x)=\frac{\hat{p}_{\delta}(x)}{\hat{\sigma}^{2}(x)}$
+is approximated by $\int_{A_{\Delta}(x)}\widehat{\sigma}_{\Delta}^{2}(x)f_{X}(t)\,dt=\widehat{\sigma}_{\Delta}^{2}(x)p_{\Delta}(x)$.
+Hence, we can estimate $\tau_{\Delta}(x)=p_{\Delta}(x)^{2}/s_{\Delta}(x)$ by
+$\widehat{\tau}_{\Delta}(x)=\frac{\widehat{p}_{\Delta}(x)}{\widehat{\sigma}_{\Delta}^{2}(x)}$
 and obtain the rate assumed in Corollary 2.6.
-The considered classical histogram estimator is a relatively simple estimator, but it has some similarities to random forests, in particular, the piecewise constant approximation.
-In a current work we apply a similar proof strategy as explained in the next section to derive confidence bands for centered purely random forests.
-Presumably, this proof strategy can also be used to obtain confidence bands for other nonparametric regression estimators.
 3 Proof strategy
-In this section we outline the proof. The estimator error can be decomposed as
-$\hat{m}(x)-m(x)=\hat{m}^{(m)}(x)-m(x)+\hat{m}^{(\varepsilon)}(x)$ (11)
-with approximation error $\hat{m}^{(m)}(x)-m(x)$ and stochastic error term $\hat{m}^{(\varepsilon)}(x)$, where
-$\hat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\qquad\text{and}$
-$\hat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
+In this section we outline the proof. The estimation error can be decomposed as
+$\widehat{m}(x)-m(x)=\widehat{m}^{(m)}(x)-m(x)+\widehat{m}^{(\varepsilon)}(x)$ (15)
+with approximation error $\widehat{m}^{(m)}(x)-m(x)$ and stochastic error $\widehat{m}^{(\varepsilon)}(x)$, where
+$\widehat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\qquad\text{and}$
+$\widehat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}.$
 First, we discuss a uniform bound on the approximation error.
-Note that $\hat{m}^{(m)}(x)=\mathbb{E}\left[\hat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise.
+Note that $\widehat{m}^{(m)}(x)=\mathbb{E}\left[\widehat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise.
 Therefore, the approximation error is indeed caused by the ability of the piecewise constant estimator to approximate $m$.
 Moreover, note that the expectation of the approximation error is the bias of the estimator.
-For any $A\subset\mathbb{R}^{p}$ let us denote its diameter
-$\mathfrak{d}(A):=\sup_{x_{1},x_{2}\in A}\|x_{1}-x_{2}\|,$
-where $\|\cdot\|$ is the Euclidean norm.
 Suppose that there exists at least one observation in $A_{\Delta}(x)$.
-Then we apply that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in[0,1]^{p}$ that
-$|\hat{m}^{(m)}(x)-m(x)|$ $=\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}$
+Then we use that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in\mathcal{X}$ that
+$|\widehat{m}^{(m)}(x)-m(x)|$ $=\Big|\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\Big|$
 $\leq C_{H}\sum_{j=1}^{n}\|X_{j}-x\|^{\alpha}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}$
-$\leq C_{H}\mathfrak{d}(A_{\delta}(x))^{\alpha}=C_{H}(\sqrt{p}\delta)^{\alpha},$ (12)
-where we used $\mathfrak{d}(A_{\delta}(x))=\delta\sqrt{p}$ by the Euclidean geometry of the cells.
-In particular, the approximation error profits from a smaller diameter or mesh size $\delta$, respectively, of the grid cells.
-To handle the stochastic error, we can replace the denominator in $\hat{m}^{(\varepsilon)}(x)$ by its expectation, such that
-$\tilde{m}^{(\varepsilon)}(x):=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}p_{\delta}(x)^{-1}\mathbb{I}\{X_{j}\in A_{\delta}(x)\}$ (13)
-becomes the asymptotically leading term of the stochastic error $\hat{m}^{(\varepsilon)}(x)$.
-Handling the remainder term $\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 4.3.
-The idea for the leading term is that $\tilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough.
+$\leq C_{H}\operatorname{diam}(A_{\Delta}(x))^{\alpha}=C_{H}C_{p}^{\alpha}\Delta^{-\alpha/p},$ (16)
+where we used $\operatorname{diam}(A_{\Delta}(x))\leq C_{p}\Delta^{-1/p}$ by Assumption 2.1. In particular, the approximation error profits from a finer partition.
+To handle the stochastic error, we can replace the denominator in $\widehat{m}^{(\varepsilon)}(x)$ by its expectation, such that
+$\widetilde{m}^{(\varepsilon)}(x)=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{p_{\Delta}(x)}$ (17)
+becomes the asymptotically leading term of the stochastic error $\widehat{m}^{(\varepsilon)}(x)$.
+Handling the remainder term $\widehat{m}^{(\varepsilon)}(x)-\widetilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 6.3.
+The idea for the leading term is that $\widetilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough.
 We write $\tau_{\Delta}(x)=p_{\Delta}(x)^{2}/s_{\Delta}(x)$, see (8) and (14).
-Note that $p_{\delta}(x)$, $\tau_{\delta}(x)$ and $s_{\delta}(x)$ only depend on the hypercube $A_{\delta}(x)$, so they take only finitely many values for $x\in[0,1]^{p}$.
+Note that $p_{\Delta}(x)$, $\tau_{\Delta}(x)$ and $s_{\Delta}(x)$ only depend on $x$ via the cell $A_{\Delta}(x)$, so they take only finitely many values for $x\in\mathcal{X}$.
 We now calculate the covariance
-$\mathrm{Cov}(\tilde{m}^{(\varepsilon)}(x_{1}),\tilde{m}^{(\varepsilon)}(x_{2}))$ $=$ $\frac{1}{n}\mathbb{E}\Big{[}\mathbb{E}\left[\varepsilon^{2}|X\right]p_{\delta}(x_{1})^{-1}p_{\delta}(x_{2})^{-1}\mathbb{I}\{X\in A_{\delta}(x_{1})\}\mathbb{I}\{X\in A_{\delta}(x_{2})\}\Big{]}$
+$\mathrm{Cov}(\widetilde{m}^{(\varepsilon)}(x_{1}),\widetilde{m}^{(\varepsilon)}(x_{2}))$ $=\frac{1}{n}\mathbb{E}\Big[\mathbb{E}\left[\varepsilon^{2}|X\right]\frac{\mathbb{I}\{X\in A_{\Delta}(x_{1})\}\mathbb{I}\{X\in A_{\Delta}(x_{2})\}}{p_{\Delta}(x_{1})p_{\Delta}(x_{2})}\Big]$
 $=\mathbb{I}\{A_{\Delta}(x_{1})=A_{\Delta}(x_{2})\}\frac{1}{n}\frac{s_{\Delta}(x_{1})}{p_{\Delta}(x_{1})^{2}}$
@@
 To obtain a process with unit variance, we consider
-$(n\tau_{\delta}(x))^{1/2}\tilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\delta}$ (14)
+$(n\tau_{\Delta}(x))^{1/2}\widetilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\Delta}$ (18)
 for the empirical process $\mathbb{G}_{n}f=\sqrt{n}(P_{n}f-Pf)$ based on independent $(X_{1},\varepsilon_{1}),\dots,(X_{n},\varepsilon_{n})$ with distribution $P$ and empirical distribution $P_{n}$, indexed by $f\in\mathcal{F}_{\Delta}$ with the finite function class
-$\mathcal{F}_{\delta}:=\big{\{}f_{x,\delta}\colon[0,1]^{p}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\delta}(t,\epsilon)=\epsilon\,s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\},x\in[0,1]^{p}\big{\}}.$ (15)
-We apply the result by Chernozhukov et al., 2014b formulated in Theorem A.1 in the Appendix A to approximate $\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\delta}$ with
+$\mathcal{F}_{\Delta}:=\big\{f_{x,\Delta}\colon\mathcal{X}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\Delta}(t,\epsilon)=\epsilon\,s_{\Delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\Delta}(x)\},x\in\mathcal{X}\big\}.$ (19)
+We apply the result by Chernozhukov et al., 2014b formulated in Theorem B.1 in the Appendix B to approximate $\sup_{f\in\mathcal{F}_{\Delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\Delta}$ with
 $\mathbf{S}_{\Delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\Delta}}|B_{\Delta}(f)|.$
@@
 $=\mathbb{I}\{A_{\Delta}(x_{1})=A_{\Delta}(x_{2})\}.$ (20)
-In particular, this implies $\mathrm{Var}(B_{\delta}(f))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables.
-The application of Theorem A.1 thus leads to the following theorem, which is proved in Section 4.
+In particular, this implies $\mathrm{Var}(B_{\Delta}(f_{x,\Delta}))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables.
+The application of Theorem B.1 thus leads to the following theorem, which is proved in Section 6.
 Theorem 3.1.
-Let $\log\delta^{-1}=\mathcal{O}(\log n)$.
-Assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$.
+Grant Assumption 2.1 and let $\log\Delta=\mathcal{O}(\log n)$.
+Assume that $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$ and $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$.
 Then there exists a sequence of random variables
-$\mathbf{S}_{\delta}\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|$
+$\mathbf{S}_{\Delta}\overset{d}{=}\max_{j=1,\ldots,\Delta}|Z_{j}|$
 with independent standard normally distributed $(Z_{j})_{j\in\mathbb{N}}$ such that
-$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right).$
+$\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right).$
 As mentioned above this direct approximation of the supremum avoids to approximate the whole empirical process uniformly by a Gaussian process.
@@
 From the covariance calculation above and the assumptions on the covariate density and variance function it follows that the variance of the stochastic error is given by
-$\mathrm{Var}\left(\tilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\delta}(x)}=\mathcal{O}\Big{(}\frac{1}{n\delta^{p}}\Big{)},$
+$\mathrm{Var}\left(\widetilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\Delta}(x)}=\mathcal{O}\Big(\frac{\Delta}{n}\Big),$
 i.e.
-the stochastic error profits from a larger $\delta$, which is caused by larger cells leading to more $\varepsilon_{j}$ being averaged and reducing the variance.
+the stochastic error profits from a partition into fewer cells as larger cells lead to more $\varepsilon_{j}$ being averaged.
 Compared to the bound of the approximation error in (16), we recover the typical bias-variance trade-off in nonparametric statistics.
 The assumptions on $\Delta$ in Theorem 2.2 lead to the commonly used undersmoothing.
-That means $\delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error allowing to build the confidence bands based on the distribution of the stochastic error.
+That means $\Delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error, allowing the confidence bands to be built based on the distribution of the stochastic error.
+4 Numerical illustration
+To illustrate our theoretical performance in a simulation example, we consider the following model: Let $p=2$, $X\sim U([0,1]^{2})$ and the regression function by given by
+$m(x)=3\sqrt{\big|0.36-\|x\|^{2}\big|}-(x_{1}+x_{2}),\qquad x=(x_{1},x_{2})\in[0,1]^{2}.$
+Note that $m$ is Hölder regular of order $\alpha\leq 1/2$ and that $m$ is close the a radial function.
+We choose homogeneous Gaussian noise with known noise level $\sigma>0$ and fix the sample size $n=800$.
+We use histogram estimators based on two different partitions.
+First, we use a histogram regression estimator $\widehat{m}_{\text{square}}$ based on a partition into $\Delta=h^{-2}$ squares of edge length $h=1/6$.
+Second, we consider an estimator $\widehat{m}_{\text{circ}}$ which relies on a partition based on circular segments.
+More precisely, we decompose the unit square into 15 circles and split the middle ones into two to four segments.
+Both partitions consist of $\Delta=36$ cells.
+For comparison we also calculate a Nadaraya-Watson estimator $\widehat{m}_{\text{nw}}$ based on a Gaussian kernel.
+The bandwidth is chosen either as $h/\sqrt{12}$, such that the standard deviation of the kernel corresponds to that of a uniform distribution on a $h\times h$ square, or via the automatic bandwidth choice implemented in the R package np.
+The true regression function, both histogram estimators as well as the Nadaraya-Watson estimator are illustrated in Figure 2.
+Figure 2: Regression function (top left) together with realizations of the Nadaraya-Watson estimator with automatic bandwidth choice (top right) and the two histogram estimators (bottom row) with two different partitions based on $n=800$ observations with noise level $\sigma=1$.
+$\sigma$ 1 2 3 4 5
+$R(\widehat{m}_{\text{circ}})$ 0.9010 1.4501 2.0383 2.6376 3.2394
+$R(\widehat{m}_{\text{square}})$ 1.1895 1.5800 2.0947 2.6338 3.1805
+$R(\widehat{m}_{\text{nw}})$ (fixed $h$) 1.0076 1.8232 2.7005 3.5900 4.4815
+$R(\widehat{m}_{\text{nw}})$ (adaptive $h$) 0.9446 1.2055 1.4417 1.6066 1.7530
+CP for $\widehat{m}_{\text{circ}}$ $c_{\Delta}(0.95)$ 0.99 1.00 1.00 1.00 1.00
+$\widetilde{c}_{\Delta}(0.95)$ 1.00 1.00 1.00 1.00 1.00
+CP for $\widehat{m}_{\text{square}}$ $c_{\Delta}(0.95)$ 0.98 0.99 0.99 1.00 1.00
+$\widetilde{c}_{\Delta}(0.95)$ 0.98 1.00 1.00 1.00 1.00
+UCP for $\widehat{m}_{\text{circ}}$ $c_{\Delta}(0.95)$ 0.05 0.61 0.78 0.85 0.88
+$\widetilde{c}_{\Delta}(0.95)$ 0.15 0.71 0.85 0.93 0.95
+UCP for $\widehat{m}_{\text{square}}$ $c_{\Delta}(0.95)$ 0.00 0.15 0.49 0.58 0.61
+$\widetilde{c}_{\Delta}(0.95)$ 0.00 0.26 0.54 0.63 0.70
+Table 1: Mean uniform estimation errors R(m^)R($\widehat${m}) for histogram estimators based on circular and quadratic partitions and mean uniform estimation error of a Nadaraya-Watson estimator; average pointwise coverage probabilities (CP) and uniform coverage probabilities (UCP) of the confidence bands with nominal level $0.95$ based on both histogram estimators with exact and asymptotic critical value, all for different noise levels $\sigma$.
+We compare the performance of the three estimators in terms of the mean uniform estimation error $R(\widehat{m})=\mathbb{E}[\sup_{x\in[0,1]^{2}}|\widehat{m}(x)-m(x)|]$ based on a grid on $0,1]^{2}$ with mesh size 0.005 and a Monte Carlo approximation of the expectation with 100 Monte Carlo iterations.
+Moreover, we calculate the uniform empirical coverage probabilities of the confidence bands with nominal confidence level $0.95$ based on the two histogram estimators and using the critical values $c_{\Delta}$ and $\widetilde{c}_{\Delta}$ from (7) and (12), respectively.
+As an additional feedback on the coverage properties of the confidence bands, we also report the average pointwise coverage probabilities over the spatial grid.
+Since the radius of the confidence bands based on $\widetilde{c}_{\Delta}$ are by the factor $\frac{\widetilde{c}_{16}(0.95)}{c_{16}(0.95)}\approx 1.0453$ larger than those based on $c_{\Delta}$, the coverage probabilities of the latter are throughout smaller.
+To see the dependence of the signal-to-noise ratio we vary the noise level $\sigma\in\{1,2,\dots,5\}$.
+All results are summarized in Table 1.
+We see that for small noise levels the performance of the histogram estimators is comparable to that of the Nadaraya-Watson estimator, while the coverage of the confidence bands is poor.
+By increasing the noise level, we enter an under-smoothing regime such that the estimation error becomes dominated by the stochastic error and the coverage considerably improves.
+Naturally, the estimation error grows with $\sigma$.
+The increasing deviation to the (adaptive) Nadaraya-Watson estimator emphasizes that the partitions are coarse in these cases.
+The partition of $\widehat{m}_{\text{circ}}$ is better adapted to the geometry of the regression function.
+As a consequence, the coverage probabilities for $\widehat{m}_{\text{circ}}$ are considerably better than for $\widehat{m}_{\text{square}}$ in all cases.
+Moreover, for noise levels $\sigma=1,2,3$ the estimation error of $\widehat{m}_{\text{circ}}$ is smaller, while for $\sigma=4,5$ the estimation error of $\widehat{m}_{\text{square}}$ is only slightly better.
+5 Conclusion
+Histogram regression estimators are used a lot in applications and serve as a building block for more advanced statistical methods such as random forests.
+Aiming for uniform statistical inference, we construct histogram based confidence bands which allow for a general choice of partitions.
+This flexibility can improve the finite sample behaviour of the histogram estimator depending on the data structure.
+Our confidence band is applicable for possibly irregular nonparametric regression functions with Hölder-regularity only of order $\alpha\in(0,1]$.
+The authors are not aware of any uniform confidence band for non-differentiable regression functions in the literature.
+These mild model assumptions are possible due to a proof structure, which deviates from more classical proof methods and which could also be transferred to other methods and models.
+For instance, the current preprint Neumeyer et al., (2025) demonstrates that a similar proof strategy can be used to derive confidence bands for centered purely random forests.
+Some ideas to obtain confidence bands for other nonparametric regression estimators are discussed in Appendix A for different choices of cells including the case of a Nadaraya-Watson estimator.
 6 Proofs
@@
 We will write $A\lesssim B$ equivalently for $A=\mathcal{O}(B)$.
-Moreover, we set $K:=\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $0,1]^{p}$.
+Moreover, we set $K=\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $\mathcal{X}$.
 6.1 Proof of Theorem 2.2 and its corollaries
 Proof of Theorem 2.2.
-We use the decomposition of the estimator and error from (11) and the definition of $\tilde{m}^{(\varepsilon)}$ in (13).
+We use the decomposition of the estimator and error from (15) and the definition of $\widetilde{m}^{(\varepsilon)}$ in (17).
 For the sequence of random variables $\mathbf{S}_{\Delta}$ as defined in Theorem 3.1 we obtain by the reverse triangle inequality that
-$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}$ $\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\mathbf{S}_{\delta}\bigg{|}\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
-$T_{1}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\hat{m}^{(m)}(x)-m(x)\big{|},$ (17)
-$T_{2}$ $=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|},$ (18)
-$T_{3}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\tilde{m}^{(\varepsilon)}(x)-\hat{m}^{(\varepsilon)}(x)|.$ (19)
+$\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}$ $\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|-\mathbf{S}_{\Delta}\bigg|\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
+$T_{1}$ $=\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}\big|\widehat{m}^{(m)}(x)-m(x)\big|,$ (21)
+$T_{2}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|,$ (22)
+$T_{3}$ $=\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}\big|\widetilde{m}^{(\varepsilon)}(x)-\widehat{m}^{(\varepsilon)}(x)|.$ (23)
 Subsequently, we will verify that all three terms are $o_{\mathbb{P}}((\log n)^{-1})$.
 For $T_{1}$ from (21) note that the approximation error satisfies
-$\hat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
+$\widehat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}$
 $\qquad+m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\Delta}(x)\}-1).$ (24)
 Using (16), we get
-$\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathfrak{d}(A_{\delta}(x))^{\alpha}\right]$
-$=C_{H}\delta^{\alpha}p^{\alpha/2}.$ (21)
-Since the number of cells is equal to $\delta^{-p}$, we obtain
-$\mathbb{E}$ $\left[\sup_{x\in[0,1]^{p}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)|\right]$
-$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\right]$
-$\leq\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}\mathbb{E}\big{[}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\big{]}$
-$=\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}(1-p_{\delta}(x))^{n}\leq\|m\|_{\infty}\delta^{-p}(1-c_{X}\delta^{p})^{n}.$
-To handle (17), note that $n^{1/2}\sup_{x}|\tau_{\delta}(x)|^{1/2}\leq(C_{X}/c_{\sigma^{2}})^{1/2}n^{1/2}\delta^{p/2}$ by (8). Together with (20) and (21) we obain for all $\kappa>0$
-$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\hat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/C_{X})^{1/2}}{n^{1/2}\delta^{p/2}\log n}\right)$
-$\lesssim\mathbb{E}\big{[}\|m-\hat{m}^{(m)}\|_{\infty}\big{]}\kappa^{-1}(n\delta^{p})^{1/2}\log n$
-$\lesssim(\delta^{2\alpha}n\delta^{p})^{1/2}\log n+\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n.$
-The first term is $o(1)$ with (3) and for the second term we have
-$\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n\lesssim\exp(-c_{X}n\delta^{p})\delta^{-p}(n\delta^{p})^{1/2}\log n\to 0$
-because $\delta\leq n^{-1/(2\alpha+p)}$ by (3) and $\delta^{p}n\geq n^{2\nu}$ by (4). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.
+$\mathbb{E}\left[\sup_{x\in\mathcal{X}}\Big|\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\Big|\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in\mathcal{X}}\operatorname{diam}(A_{\Delta}(x))^{\alpha}\right]$
+$=C_{H}C_{p}^{\alpha}\Delta^{-\alpha/p}.$ (25)
+Moreover, we obtain by (10)
+$\mathbb{E}$ $\left[\sup_{x\in\mathcal{X}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\Delta}(x)\}-1)|\right]$
+$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in\mathcal{X}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\Delta}(x)\}\right]$
+$\leq\|m\|_{\infty}\Delta\sup_{x\in\mathcal{X}}\mathbb{E}\big[\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\Delta}(x)\}\big]$
+$=\|m\|_{\infty}\Delta\sup_{x\in\mathcal{X}}(1-p_{\Delta}(x))^{n}\leq\|m\|_{\infty}\Delta(1-c_{X}{c_{p}\Delta^{-1}})^{n}.$
+To handle (21), note that $n^{1/2}\sup_{x}|\tau_{\Delta}(x)|^{1/2}\leq(C_{X}C_{p}/c_{\sigma^{2}})^{1/2}(n/\Delta)^{1/2}$ by (11). Together with (24) and (25) we obtain for all $\kappa>0$
+$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\widehat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/(C_{X})C_{p})^{1/2}\Delta^{1/2}}{n^{1/2}\log n}\right)$
+$\lesssim\mathbb{E}\big[\|m-\widehat{m}^{(m)}\|_{\infty}\big]\kappa^{-1}(n/\Delta)^{1/2}\log n$
+$\lesssim{(\Delta^{-2\alpha/p}n/\Delta)^{1/2}\log n+\Delta(1-c_{X}c_{p}\Delta^{-1})^{n}(n/\Delta)^{1/2}\log n}.$
+The first term is $o(1)$ due to (5) and for the second term we have
+$\Delta(1-c_{X}c_{p}\Delta^{-1})^{n}(n/\Delta)^{1/2}\log n\lesssim\exp(-c_{X}c_{p}n/\Delta)\Delta(n/\Delta)^{1/2}\log n\to 0$
+because $\Delta\geq n^{1/(2\alpha/p+1)}$ by (5) and $n/\Delta\geq n^{2\nu}$ by (6). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.
 For the term $T_{2}$ from (22) Theorem 3.1 yields
-$T_{2}=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$ (22)
+$T_{2}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|$
+$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right)$ (26)
 because $\nu\geq 4$. To see that the right-hand side is indeed $o_{\mathbb{P}}((\log n)^{-1})$, note that assumption (6) implies
-$\frac{(\log n)^{5/2}}{\delta^{p/2}n^{1/2-1/\nu}}=\Big{(}\frac{(\log n)^{5}}{\delta^{p}n}n^{2/\nu}\Big{)}^{1/2}\to 0.$
+$\frac{(\log n)^{5/2}\Delta^{1/2}}{n^{1/2-1/\nu}}=\Big(\frac{(\log n)^{5}\Delta}{n}n^{2/\nu}\Big)^{1/2}\to 0.$
 The same argument applied to the other terms in (26) yields $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$.
-Next we consider the remainder term $T_{3}$ from (19). By the definition of $\tilde{m}^{(\varepsilon)}$ in (13) and $\hat{p}_{\delta}(x)$ in (9) we obtain
-$\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)=\tilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right).$
+Next we consider the remainder term $T_{3}$ from (23). By the definition of $\widetilde{m}^{(\varepsilon)}$ in (17) and $\widehat{p}_{\Delta}(x)$ in (13) we obtain
+$\widehat{m}^{(\varepsilon)}(x)-\widetilde{m}^{(\varepsilon)}(x)=\widetilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\Delta}(x)}{\widehat{p}_{\Delta}(x)}\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-1\right).$
 Together with (9) and $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$, we conclude
-$T_{3}$ $\leq$ $\left(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in[0,1]^{p}}\left|\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right|$
-$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-\mathbb{I}\{\hat{p}_{\delta}(x)=0\}\right|.$
+$T_{3}$ $\leq$ $\left(\mathbf{S}_{\Delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in\mathcal{X}}\left|\frac{p_{\Delta}(x)}{\widehat{p}_{\Delta}(x)}\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-1\right|$
+$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in\mathcal{X}}\left|np_{\Delta}(x)\left(\frac{1}{n\widehat{p}_{\Delta}(x)}-\frac{1}{np_{\Delta}(x)}\right)\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}\right|.$
 To obtain the remainder rate $o_{\mathbb{P}}((\log n)^{-1})$, the second indicator can be written as $(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\Delta}(x)\}-1)$ and be handled as in (24).
-For the other term note that for fixed $x$ the random variable $B_{n}=n\hat{p}_{\delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\delta}(x)$, and we can apply Lemma 4.1.
+For the other term note that for fixed $x$ the random variable $B_{n}=n\widehat{p}_{\Delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\Delta}(x)$, and we can apply Lemma 6.1.
 Together with (10), the union bound and Markov’s inequality we obtain
-$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\right|\mathbb{I}\{\hat{p}_{\delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
-$\leq$ $\delta^{-p}\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}\delta^{p})^{q}\mathcal{O}((n\delta^{p})^{-3q/2})=\mathcal{O}\left(\delta^{-p}\frac{(\log n)^{3q/2}}{(\delta^{p}n)^{q/2}}\right)=o(1)$
+$\mathbb{P}\left(\sup_{x\in\mathcal{X}}\left|np_{\Delta}(x)\left(\frac{1}{n\widehat{p}_{\Delta}(x)}-\frac{1}{np_{\Delta}(x)}\right)\right|\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
+$\leq\Delta\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}C_{p}/\Delta)^{q}\mathcal{O}((n/\Delta)^{-3q/2})=\mathcal{O}\left(\Delta\frac{(\log n)^{3q/2}}{(n/\Delta)^{q/2}}\right)=o(1)$ (27)
 which holds by assumption (6) for $q\in 2\mathbb{N}$, $q\geq\nu$. Hence, $T_{3}=o_{\mathbb{P}}((\log n)^{-1})$.
-Altogether we have now proved
-$\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|=\mathbf{S}_{\delta}+R_{n}$ (23)
+Altogether we conclude
+$\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|=\mathbf{S}_{\Delta}+R_{n}$ (28)
 for a remainder term $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$. Thus, the uniform coverage probability can be bounded from below by
-$\mathbb{P}$ $\big{(}m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\big{)}$
-$=\mathbb{P}\Big{(}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\leq c_{\delta}(\beta)\Big{)}$
-$=\mathbb{P}\big{(}\mathbf{S}_{\delta}+R_{n}\leq c_{\delta}(\beta)\big{)}$
+$\mathbb{P}$ $\big(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\big)$
+$=\mathbb{P}\Big(\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\leq c_{\Delta}(\beta)\Big)$
+$=\mathbb{P}\big(\mathbf{S}_{\Delta}+R_{n}\leq c_{\Delta}(\beta)\big)$ (29)
 $\geq\mathbb{P}\big(\mathbf{S}_{\Delta}\leq c_{\Delta}(\beta)\big)-\mathbb{P}\big(c_{\Delta}(\beta)-\kappa_{n}<\mathbf{S}_{\Delta}\leq c_{\Delta}(\beta)\big)-\mathbb{P}\left(R_{n}>\kappa_{n}\right),$
@@
 $\sup_{\xi\in\mathbb{R}}\mathbb{P}\left(|\mathbf{S}_{\Delta}-\xi|\leq\kappa_{n}\right)$ $\leq 4\kappa_{n}\left(\mathbb{E}\left[\sup_{f\in\mathcal{F}_{\Delta}}|B_{\Delta}(f)|\right]+1\right)$
-$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\right]+1\right)$
-$=\kappa_{n}\mathcal{O}(\sqrt{\log n})=o(1)$
+$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\Delta}|Z_{j}|\right]+1\right)$
+$=\mathcal{O}(\kappa_{n}\sqrt{\log n})=o(1)$
 which follows from Corollary 2.1 by Chernozhukov et al., 2014a and (9). Thus we obtain
-$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.\qed$
+$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.\qed$
+Proof of Corollary 2.4.
+As in (28) and (29) we obtain
+$\mathbb{P}\big(m(x)\in\widetilde{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\big)$ $=$ $\mathbb{P}\big(\mathbf{S}_{\Delta}+R_{n}\leq\widetilde{c}_{\Delta}(\beta)\big)$
+$=$ $\mathbb{P}\big(a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})+a_{\Delta}R_{n}\leq-\log(-\log(1-\beta))\big)$
+$=$ $\mathbb{P}\big(a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})\leq-\log(-\log(1-\beta))\big)+o(1).$
+The last equality follows from Slutsky’s lemma because $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$, and from assumption (6) it follows that $a_{\Delta}\lesssim(\log n)^{1/2}$.
+It remains to verify that the limit distribution of $a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})\overset{d}{=}a_{\Delta}(\max_{j=1,\ldots,\Delta}|Z_{j}|-b_{\Delta})$ is the Gumbel distribution with cdf $\exp(-e^{-x})$.
+Indeed, using the the asymptotic independence of the maximum and minimum (Leadbetter et al.,, 2012, Theorem 1.8.2) together with the extreme value limit of the Gaussian distribution (Leadbetter et al.,, 2012, Theorem 1.5.3) we have
+$\mathbb{P}(a_{\Delta}(\max_{j}|Z_{j}|-b^{\prime}_{\Delta}$ $)\leq x)=\mathbb{P}\Big(\max_{j}Z_{j}\leq\frac{x}{a_{\Delta}}+b^{\prime}_{\Delta},\min_{j}Z_{j}\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)$
+$=\mathbb{P}\Big(\max_{j}Z_{j}\leq\frac{x}{a_{\Delta}}+b^{\prime}_{\Delta}\Big)\mathbb{P}\Big(\min_{j}Z_{j}\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)+o(1)$
+$=\mathbb{P}\Big(a_{\Delta}(\max_{j}Z_{j}-b^{\prime}_{\Delta})\leq x\Big)\mathbb{P}\Big(-\max_{j}(-Z_{j})\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)+o(1)$
+$=\mathbb{P}\Big(a_{\Delta}(\max_{j}Z_{j}-b^{\prime}_{\Delta})\leq x\Big)^{2}+o(1)\to e^{-2e^{-x}}$
+for $a_{\Delta}=\sqrt{2\log\Delta}$ and $b^{\prime}_{\Delta}=\sqrt{2\log\Delta}-\frac{\log\log n+\log 4\pi}{2\sqrt{2\log\Delta}}$.
+Hence, $b_{\Delta}=b^{\prime}_{\Delta}+a_{\Delta}^{-1}\log 2$ yields $\mathbb{P}(a_{\Delta}(\max|Z_{j}|-b_{\Delta})\leq x)\to e^{-e^{-x}}$.
+∎
 Proof of Corollary 2.5.
@@
 The estimates following (24) and (11) show that the approximation error satisfies
-$\mathbb{E}\big{[}\sup_{x\in[0,1]^{p}}|\hat{m}^{(m)}(x)-m(x)|\big{]}\lesssim n^{-1/2}\delta^{-p/2}\mathbb{E}[T_{1}]\lesssim\delta^{\alpha}+\delta^{-p}(1-c_{X}\delta^{p})^{n},$
-where the second term is negligible if $\log(\delta^{-1})=o(n\delta^{p})$.
+$\mathbb{E}\big[\sup_{x\in\mathcal{X}}|\widehat{m}^{(m)}(x)-m(x)|\big]\lesssim n^{-1/2}\Delta^{1/2}\mathbb{E}[T_{1}]\lesssim\Delta^{-\alpha/p}+\Delta(1-c_{X}c_{p}/\Delta)^{n},$
+where the second term is negligible if $\log(\Delta)=o(n/\Delta)$.
 The latter condition is implied by assumption (6).
-Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (6) and (8), we conclude $\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}})$.
-Choosing $\delta=(n/\log n)^{-1/(2\alpha+p)}$ balances both terms and yields the optimal rate of convergence.
+Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (9) and (11), we conclude $\sup_{x\in\mathcal{X}}|\widehat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\Delta^{-\alpha/p}+\sqrt{(\log n)\Delta/n})$.
+Choosing $\Delta=(n/\log n)^{1/(2\alpha/p+1)}$ balances both terms and yields the optimal rate of convergence.
 ∎
@@
 As in the proof of Theorem 2.2 we need to show $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$ for the additional remainder term
-$R_{n}$ $=$ $\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in[0,1]^{p}}\hat{\tau}_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\bigg{|}$
-$\leq$ $\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)^{1/2}-\tau_{\delta}(x)^{1/2}|}{\tau_{\delta}(x)^{1/2}}$
-$=$ $(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1}))\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)-\tau_{\delta}(x)|}{\tau_{\delta}(x)^{1/2}(\tau_{\delta}(x)^{1/2}+\hat{\tau}_{\delta}(x)^{1/2})}$
-$\leq$ $\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}((\log n)^{-1})$
-by (23), (25), (6) and the assumption on $\hat{\tau}_{\delta}$. ∎
+$R_{n}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in\mathcal{X}}\widehat{\tau}_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\bigg|$
+$\leq\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\sup_{x\in\mathcal{X}}\frac{|\widehat{\tau}_{\Delta}(x)^{1/2}-\tau_{\Delta}(x)^{1/2}|}{\tau_{\Delta}(x)^{1/2}}$
+$=\big(\mathbf{S}_{\Delta}+o_{\mathbb{P}}((\log n)^{-1})\big)\sup_{x\in\mathcal{X}}\frac{|\widehat{\tau}_{\Delta}(x)-\tau_{\Delta}(x)|}{\tau_{\Delta}(x)^{1/2}(\tau_{\Delta}(x)^{1/2}+\widehat{\tau}_{\Delta}(x)^{1/2})}$
+$\leq\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{\tau}_{\Delta}(x)}{\tau_{\Delta}(x)}-1\Big|=o_{\mathbb{P}}((\log n)^{-1})$
+by (28), (32), (9) and the assumption on $\widehat{\tau}_{\Delta}$. ∎
 6.2 Proof of Theorem 3.1
 Proof of Theorem 3.1.
-We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem A.1 in the appendix to fix the notation.
+We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem B.1 in the appendix to fix the notation.
 While the function class in this result does not depend on $n$ and $\Delta$, the theorem can be applied to $\mathcal{F}_{\Delta}$ for every $n$ or $\Delta$, respectively, due to the universality of the occuring constants.
@@
 To see that $\mathcal{F}_{\Delta}$ is a VC-type class as defined in the appendix, we calculate a measurable envelope $F$:
-$\sup_{x\in[0,1]^{p}}|f_{x,\delta}(t,\epsilon)|$ $=\sup_{x\in[0,1]^{p}}|\epsilon|s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\}$
-$\leq|\epsilon|(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}=:F(t,\epsilon)$ (24)
-applying $c_{\sigma^{2}}c_{X}\delta^{p}\leq s_{\delta}(x)\leq C_{\sigma^{2}}C_{X}\delta^{p}$ from (10), (1) and (2).
-The finite size $|\mathcal{F}_{\delta}|=\delta^{-p}$ is an upper bound for any covering number of $\mathcal{F}_{\delta}$.
+$\sup_{x\in\mathcal{X}}|f_{x,\Delta}(t,\epsilon)|$ $=\sup_{x\in\mathcal{X}}|\epsilon|s_{\Delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\Delta}(x)\}$
+$\leq|\epsilon|(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}\Delta^{1/2}=:F(t,\epsilon)$ (30)
+applying $c_{\sigma^{2}}c_{X}c_{p}\Delta^{-1}\leq s_{\Delta}(x)\leq C_{\sigma^{2}}C_{X}C_{p}\Delta^{-1}$ from (14), (1), (2) and Assumption 2.1.
+The finite size $|\mathcal{F}_{\Delta}|=\Delta$ is an upper bound for any covering number of $\mathcal{F}_{\Delta}$.
 Hence,
-$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\delta^{-p}\leq\delta^{-p}/\kappa$
-for all $\kappa\in(0,1]$, $\mathcal{F}_{\delta}$ is indeed of VC-type and we can choose $A=\delta^{-p}$ and $v=1$ in Theorem A.1.
+$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\Delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\Delta\leq\Delta/\kappa$ (31)
+for all $\kappa\in(0,1]$, $\mathcal{F}_{\Delta}$ is indeed of VC-type and we can choose $A=\Delta$ and $v=1$ in Theorem B.1.
 We proceed by verifying the conditions on the moments. Since
-$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{2}=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}\right]=1,$
-$\tilde{\sigma}$ from Theorem A.1 is equal to $1$. Moreover,
-$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{3}$ $=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}|^{3}\right]$
-$\leq K(c_{\sigma^{2}}c_{X})^{-3/2}C_{X}\delta^{-p/2}.$
+$\sup_{f\in\mathcal{F}_{\Delta}}P|f|^{2}=\sup_{x\in\mathcal{X}}s_{\Delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\Delta}(x)\}\right]=1,$
+the quantity $\widetilde{\sigma}$ from Theorem B.1 is equal to $1$. Moreover,
+$\sup_{f\in\mathcal{F}_{\Delta}}P|f|^{3}$ $=\sup_{x\in\mathcal{X}}s_{\Delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\Delta}(x)\}|^{3}\right]$
+$\leq K(c_{\sigma^{2}}c_{X}c_{p})^{-3/2}C_{X}C_{p}\Delta^{1/2}.$
 For the envelope $F$ from (30) we obtain
-$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
-We thus choose $b$ from Theorem A.1 as
-$b:=\delta^{-p/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}})^{-3/2}C_{X},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}})^{-1/2}\right\}.$
-Since $b$ is of order $\delta^{-p/2}$, we have $b\geq 1$ for $\delta$ small enough. To apply Theorem A.1 we obtain
-$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\tilde{\sigma}))=c(\log n\vee\log(\delta^{-p}\delta^{-p/2}C_{b}))$
-$\lesssim\log n\vee p\log(\delta^{-1})\lesssim\log n.$
-Let $B_{\delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem A.1 yields that there exists a random variable
-$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}}B_{\delta}(f)=\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|=\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|$
+$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}\Delta^{1/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
+We thus choose $b$ from Theorem B.1 as
+$b:=\Delta^{1/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}}C_{p})^{-3/2}C_{X}C_{p},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}}C_{p})^{-1/2}\right\}.$
+Since $b$ is of order $\Delta^{1/2}$, we have $b\geq 1$ for $\Delta$ large enough. To apply Theorem B.1 we obtain
+$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\widetilde{\sigma}))=c(\log n\vee\log(\Delta^{3/2}C_{b}))$
+$\lesssim\log n\vee p\log(\Delta)\lesssim\log n.$
+Let $B_{\Delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem B.1 yields that there exists a random variable
+$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\Delta}\cup-\mathcal{F}_{\Delta}}B_{\Delta}(f)=\sup_{f\in\mathcal{F}_{\Delta}}|B_{\Delta}(f)|=\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|$
 such that
@@
 In view of (18), we conclude
-$\bigg{|}$ $\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}$
+$\bigg|$ $\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|$
 $=\Big|\sup_{f\in\mathcal{F}_{\Delta}}|\mathbb{G}_{n}f|-\mathbf{S}_{\Delta}\Big|$
 $=\mathcal{O}_{\mathbb{P}}\left(\frac{b\Gamma_{n}(\log n)^{1/2}}{n^{1/2-1/\nu}}+\frac{(b\log n)^{1/2}\Gamma_{n}^{3/4}}{n^{1/4}}+\frac{(b\Gamma_{n}^{2}\log n)^{1/3}}{n^{1/6}}\right)$
-$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$
-since $\Gamma_{n}=\mathcal{O}(\log n)$ and $b=\mathcal{O}(\delta^{-p/2})$. Finally note, that the structure of the covariance (16) implies that
-$\mathbf{S}_{\delta}\overset{d}{=}\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|,$ (25)
+$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right)$
+since $\Gamma_{n}\lesssim\log n$ and $b\lesssim\Delta^{1/2}$. Finally note, that the structure of the covariance (20) implies that
+$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|\overset{d}{=}\max_{j=1,\ldots,\Delta}|Z_{j}|,$ (32)
 for independent standard normally distributed $Z_{1},Z_{2},\dots$. ∎
@@
 Lemma 6.2.
-If (4) is satisfied, we have
-$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}\big{(}(\log n)^{-3/2})\big{)}.$
+Let Assumption 2.1 and (6) be satisfied. Then we have
+$\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|=o_{\mathbb{P}}\big((\log n)^{-3/2})\big).$
 Proof.
 Note that
-$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}$ $=$ $\sup_{x\in[0,1]^{p}}\frac{1}{np_{\delta}(x)}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}.$
-Since $x\mapsto\hat{p}_{\delta}(x)$ is piecewise constant with at most $\delta{-p}$ different values, the supremum can be handled with a union bound.
+$\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|$ $=\sup_{x\in\mathcal{X}}\frac{1}{np_{\Delta}(x)}\Big|\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}-np_{\Delta}(x)\Big|.$
+Since $x\mapsto\widehat{p}_{\Delta}(x)$ is piecewise constant with at most $\Delta$ different values, the supremum can be handled with a union bound.
 Moreover, we can exploit that $\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}$ is $Bin(n,p_{\Delta}(x))$-distributed.
 The binomial moment result (33) and (10) yield for any $q\in\mathbb{N}$ that
-$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
-$\qquad\leq\delta^{-p}\frac{1}{(nc_{X}\delta^{p})^{q}}\mathbb{E}\Big{[}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}^{q}\Big{]}\frac{(\log n)^{3q/2}}{\kappa^{q}}$
-$\qquad\lesssim\frac{\delta^{-p}(\log n)^{3q/2}}{\kappa^{q}(\delta^{p}n)^{q/2}}=o(1)$
+$\mathbb{P}\left(\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
+$\qquad\leq\frac{\Delta}{(nc_{X}c_{p}/\Delta)^{q}}\mathbb{E}\Big[\Big|\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}-np_{\Delta}(x)\Big|^{q}\Big]\frac{(\log n)^{3q/2}}{\kappa^{q}}$
+$\qquad\lesssim\frac{\Delta(\log n)^{3q/2}}{\kappa^{q}(n/\Delta)^{q/2}}=o(1)$
 which holds by assumption (6) for large enough $q$. ∎
-Appendix A Approximation of suprema of empirical processes
+Appendix A Extensions
+For the general partition in Assumption 2.1 we chose disjoint cells, but intersecting cells are also possible.
+For this we fix a mapping $x\mapsto A_{\Delta}(x)$, where all $(A_{\Delta}(x))_{x\in\mathcal{X}}$ satisfy (3), and we assume that $\{A_{\Delta}(x)\mid x\in\mathcal{X}\}$ is a subset of some VC-class of sets, not depending on $\Delta$.
+For instance the cells $A_{\Delta}(x)$ can be chosen as balls (or hypercubes) with middle point $x$ and radius (side length) depending on $\Delta$.
+We use the same notation for the estimator $\widehat{m}$ in (4) and the assumptions in Theorem 2.2.
+However, $\Delta$ is not the number of cells anymore, and we do not assume that there are only finitely many cells for each $\Delta$.
+Thus the parts of the proof applying the union bound and the finite cardinality of the function class $\mathcal{F}_{\Delta}$ have to be modified.
+The general proof structure is still the same as in Sections 3 and 6.
+One starts with the same expansion with terms $T_{1}$, $T_{2}$ and $T_{3}$ from (21), (22) and (23), but with an adjusted $\mathbf{S}_{\Delta}$ discussed below.
+To handle $T_{1}$, (25) works out the same way, but for (24) and showing negligibility of $T_{3}$ a different argument is needed.
+Let $P$ denote the distribution of $X_{1}$ and $P_{n}$ the empirical distribution based on $X_{1},\dots,X_{n}$.
+Then one can write
+$\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|=\sup_{g\in\mathcal{G}_{n}}|P_{n}g-Pg|$
+for the function class $\mathcal{G}_{n}=\{\mathbb{I}_{A_{\Delta}(x)}\mid x\in\mathcal{X}\}$ (with $\Delta=\Delta_{n}$), which is a subset of a VC-class not depending on $n$.
+We have $|g|\leq 1$ and $(Pg^{2})^{1/2}=(p_{\Delta}(x))^{1/2}\leq(C_{X}C_{p})^{1/2}\Delta^{-1/2}$ for all $g\in\mathcal{G}_{n}$ by (10) and (3).
+Applying Theorem 37 in Pollard, (1984, p.
+34), one obtains
+$\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|=o_{\mathbb{P}}(\alpha_{n}\Delta^{-1})$
+for each positive non-increasing sequence $\alpha_{n}$ with $(\Delta\log n)/(n\alpha_{n}^{2})=o(1)$.
+In view of (6) we can choose $\alpha_{n}=(\log n)^{-2}$.
+With this one also obtains a similar result as (27) because
+$\sup_{x\in\mathcal{X}}\left|p_{\Delta}(x)\left(\frac{1}{\widehat{p}_{\Delta}(x)}-\frac{1}{p_{\Delta}(x)}\right)\right|\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}$
+$\leq$ $\frac{\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|}{\inf_{x\in\mathcal{X}}p_{\Delta}(x)}\sup_{x\in\mathcal{X}}\frac{p_{\Delta}(x)}{p_{\Delta}(x)-\sup_{t\in\mathcal{X}}|\widehat{p}_{\Delta}(t)-p_{\Delta}(t)|}=o_{\mathbb{P}}((\log n)^{-2}).$
+Further for the term $\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}$ in the upper bound of $T_{3}$ note that for $\eta>0$
+$\mathbb{P}\left((\log n)^{3/2}\sup_{x\in\mathcal{X}}\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}>\eta\right)\leq$ $\mathbb{P}\left(\exists x\in\mathcal{X}:\widehat{p}_{\Delta}(x)=0\right)$
+$\leq$ $\mathbb{P}\left(\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|\geq c_{X}c_{p}\Delta^{-1}\right)=o(1),$
+and to handle term (24) for $T_{1}$ works in the same way.
+Thus $T_{1}$ is negligible.
+For negligibility of $T_{3}$ one also needs the rate $\mathbf{S}_{\Delta}=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$, and arguments for this are based on the cell structures.
+We consider this in the examples below.
+The dominating term $T_{2}$ can be treated as in Theorem 3.1.
+For this let $\mathcal{A}_{\Delta}(x)$ be the VC-class containing all sets $A_{\Delta}(x)$.
+Then $\mathcal{F}_{\Delta}$ with the analogous definition as in (19) and envelope $F$ from (30) is a subset of the VC-type class
+$\{(t,\epsilon)\mapsto|\epsilon|c\Delta^{1/2}\mathbb{I}_{A}\mid A\in\mathcal{A}_{\Delta}(x),c\in[0,(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}],x\in\mathcal{X}\}$
+such that the covering number is bounded by
+$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\Delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq C\kappa^{-V}$
+for some $C$ and $V$ not depending on $\Delta$.
+So we even get a better upper bound than in (31) and the remaining part of the proof of Theorem 3.1 can be done analogously.
+The approximation has the structure
+$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|,$
+where the centered Gaussian process $B_{\Delta}$ has the covariance formula
+$\mathrm{Cov}(B_{\Delta}(f_{x_{1},\Delta}),B_{\Delta}(f_{x_{2},\Delta}))=\frac{\int_{A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2})}\sigma^{2}(t)f_{X}(t)\,dt}{(s_{\Delta}(x_{1})s_{\Delta}(x_{2}))^{1/2}},$
+and the same distribution as the process
+$\frac{1}{(s_{\Delta}(x))^{1/2}}\int\sigma(t)(f_{X}(t))^{1/2}\mathbb{I}_{A_{\Delta}(x)}(t)\,dW(t)$ (35)
+with a $p$-parameter Brownian sheet $W$.
+Due to the overlapping cells, we don’t have an diagonal covariance structure anymore and $\mathbf{S}_{\Delta}$ cannot be represented by a finite number of Gaussian random variables anymore.
+Example A.1.
+Consider the homoscedastic case with $X\sim U[0,1]^{p}$. Then the covariance structure simplifies to
+$\mathrm{Cov}(B_{\Delta}(f_{x_{1},\Delta}),B_{\Delta}(f_{x_{2},\Delta}))=\frac{\operatorname{vol}(A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2}))}{(\operatorname{vol}(A_{\Delta}(x_{1}))\operatorname{vol}(A_{\Delta}(x_{2})))^{1/2}},$
+and quantiles of $\mathbf{S}_{\Delta}$ can be directly simulated to obtain a confidence band.
+For the proof discussed above one needs the convergence rate of $\mathbf{S}_{\Delta}$.
+Let us consider the case where $A_{\Delta}(x)$ is a ball around $x$ (intersected with $0,1]^{p}$) with radius $r_{\Delta}(x)$, which may depend on $x$ with the assumption
+$c_{r}\Delta^{-1/p}\leq r_{\Delta}(\cdot)\leq C_{r}\Delta^{-1/p}$
+for some constants $C_{r}\geq c_{r}>0$. Define the distance between $x_{1},x_{2}\in[0,1]^{p}$ as
+$d(x_{1},x_{2})=\big(\mathrm{Var}(B_{\Delta}(f_{x_{1},\Delta})-B_{\Delta}(f_{x_{2},\Delta}))\big)^{1/2}=\left(1-\frac{\operatorname{vol}(A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2}))}{(\operatorname{vol}(A_{\Delta}(x_{1}))\operatorname{vol}(A_{\Delta}(x_{2})))^{1/2}}\right)^{1/2},$
+and assume it can be upper bounded by
+$d(x_{1},x_{2})\leq\left(c\|x_{1}-x_{2}\|\Delta^{1/p}\right)^{1/2}$
+for some constant $c$.
+For instance for dimension $p=3$ one can apply a sphere-sphere intersection formula and assume Lipschitz continuity of the map $x\mapsto r_{\Delta}(x)$.
+Then one obtains an upper bound for the covering number $N(\epsilon,\mathcal{F}_{\Delta},d)\lesssim(c\Delta^{1/p}/\epsilon^{2})^{p}$.
+By Massart, (2007, Theorem 3.18) we obtain
+$\mathbb{E}[\mathbf{S}_{\Delta}]\lesssim\int_{0}^{1}\sqrt{\log((c\Delta^{1/p}/\epsilon^{2})^{p})}\,d\epsilon=\mathcal{O}(\sqrt{\log\Delta})$
+and thus $\mathbf{S}_{\Delta}=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$.
+Example A.2.
+Consider the special case where $t\in A_{\Delta}(x)$ is equivalent to $(x-t)/\delta\in A$ for one fixed set $A$ such that one can write $\mathbb{I}\{t\in A_{\Delta}(x)\}=K(\frac{x-t}{\delta})$ with $K=\mathbb{I}_{A}$ and for example $A=[-1,1]^{p}$.
+For the assumptions on $\Delta$ as in Example 2.3 let $\Delta=\delta^{-p}$.
+The estimator $\widehat{m}$ in (4) is then a Nadaraya-Watson estimator with kernel $K$ and bandwidth $\delta$.
+As in the last steps of classical confidence band proof structures, see Chao et al., (2017), one can now replace $\sigma(t)(f_{X}(t))^{1/2}$ inside the integral (35) by $\sigma(x)(f_{X}(x))^{1/2}$ and use a factor function to obtain the simple process
+$Z_{\delta}(x)=\frac{1}{\sqrt{\delta^{p}}}\int K\left(\frac{x-t}{\delta}\right)\,dW(t).$
+For the replacement some assumptions on the functions $\sigma$ and $f_{X}$ are needed.
+For the proof structure considered above note that instead of the rate of $\mathbf{S}_{\Delta}$ one can use the rate $\sup_{x}|Z_{\delta}(x)|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$, and one obtains the limit distribution with the classical Rosenblatt, (1976) result.
+We are not aware of literature about asymptotic confidence bands based on Nadaraya-Watson estimators in the multidimensional covariate case for regression functions with Hölder regularity $\alpha\in(0,1]$.
+This additionally demonstrates the usefulness of the new proof structure.
+Appendix B Approximation of suprema of empirical processes
 Denote by $P_{n}$ the empirical distribution of independent random variables $Z_{1},\dots,Z_{n}$ with distribution $P$.
@@
 The next theorem is Corollary 2.2 by Chernozhukov et al., 2014b .
-Theorem A.1.
+Theorem B.1.
 Suppose that $\mathcal{F}$ is a pointwise measurable function class with envelope $F$ that is VC-type for constants $A\geq e,v\geq 1$.
@@
 $\mathbb{E}[G_{P}(f)G_{P}(g)]=P(fg)=\mathbb{E}[f(Z_{1})g(Z_{1})].$
-Suppose also that for some $b\geq\tilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\tilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$.
+Suppose also that for some $b\geq\widetilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\widetilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$.
 Then for every $\gamma\in(0,1)$ there exists a random variable $\mathbf{S}\overset{d}{=}\sup_{f\in\mathcal{F}}G_{P}f$ such that
-$\mathbb{P}\Big{(}\Big{|}\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big{|}>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\tilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\tilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big{)}\leq C\Big{(}\gamma+\frac{\log n}{n}\Big{)}$
-where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\tilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).
+$\mathbb{P}\Big(\Big|\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big|>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\widetilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\widetilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big)\leq C\Big(\gamma+\frac{\log n}{n}\Big)$
+where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\widetilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).
+Acknowledgments.
+The authors are grateful to the Editor-in-Chief, Dietrich von Rosen, an Associate Editor and the reviewers for their valuable comments, which led to an improved manuscript.
+Moreover, we thank Max Farrell for his insightful remarks.